### v2

# : Generalizing Large Language Models for E-commerce

Abstract
With tremendous efforts in developing effective e-commerce models, conventional e-commerce models show limited success in generalist e-commerce modeling, and suffer from unsatisfactory performance on new users and new products β a typical out-of-domain generalization challenge. Meanwhile, large language models (LLMs) demonstrate outstanding performance in generalist modeling and out-of-domain generalizability in many fields. Toward fully unleashing their power for e-commerce, in this paper, we construct , the first open-sourced, large-scale, and high-quality benchmark instruction dataset for e-commerce. Leveraging , we develop , a series of e-commerce LLMs, by instruction-tuning general-purpose LLMs. Our comprehensive experiments and evaluation demonstrate that models substantially outperform baseline models, including the most advanced GPT-4, and the state-of-the-art task-specific models in in-domain evaluation. Moreover, exhibits excellent generalizability to out-of-domain settings, including unseen products and unseen instructions, highlighting its superiority as a generalist e-commerce model. Both the dataset and the models show great potential in empowering versatile and effective LLMs for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM/.
1 Introduction
The Internetβs evolution and the rise of the digital economy have made e-commerce an integral part of daily life, drawing considerable attention from researchers. In recent years, tremendous research efforts have been dedicated to developing effective e-commerce modelsΒ (Yang etΒ al., 2022; Geng etΒ al., 2022). Though promising, conventional e-commerce models generally suffer from two issues. (1) Limited success in generalist e-commerce modelingΒ (Zhang etΒ al., 2023b): Conventional e-commerce models are typically task-specific (e.g., gSASRec for sequential recommendationΒ (Petrov & Macdonald, 2023), SUOpenTag for attribute value extractionΒ (Xu etΒ al., 2019)). However, contemporary e-commerce platforms, such as Amazon and eBay, are highly complex and consistently expanding, with many interdependent tasks. In this context, generalist modeling is particularly suitable and desired on these platforms due to its cost-effectiveness and extensibility, while the task-specific modeling scheme struggles with scalability. (2) Unsatisfactory performance on new users and new productsΒ (Hou etΒ al., 2024): Cold start (i.e., performing e-commerce tasks on new users or new products)Β (Lika etΒ al., 2014), the typical out-of-domain (OOD) generalization challenge in e-commerce, has been a long-standing and difficult problemΒ (Ding etΒ al., 2022; Yang etΒ al., 2022). Existing task-specific e-commerce models are typically tailored to existing users and products and lack the ability to effectively extrapolate to new users and new productsΒ (Lika etΒ al., 2014). However, new users and new products are very common and highly wanted in the dynamic landscape of e-commerce.
Recently, large language models (LLMs), such as GPT-4Β (OpenAI, 2023), Claude 2Β (Anthropic, 2023), GeminiΒ (Team etΒ al., 2023), and Llama 2Β (Touvron etΒ al., 2023b), have demonstrated exceptional performance in natural language processingΒ (Zhao etΒ al., 2023), information retrievalΒ (Spatharioti etΒ al., 2023), and many other fieldsΒ (Frieder etΒ al., 2023; Zeng etΒ al., 2023). However, their power for e-commerce is not fully unleashed. While limited effortsΒ (Yang etΒ al., 2023; Zhang etΒ al., 2023a; Li etΒ al., 2024) have been dedicated to leveraging LLMs for e-commerce, most of them focus on studying the utility of pre-trained LLMs in single or homogenous tasks (e.g., e-commerce authoring). Some recent efforts employ instruction tuning to adapt LLMs for e-commerce applicationsΒ (Shi etΒ al., 2023; Li etΒ al., 2024; Geng etΒ al., 2022). However, they fall short in the covered e-commerce tasks and instruction data. We aim to bridge the gap and develop e-commerce foundation models with real-world utilities for a large variety of e-commerce applications.
To this end, we construct , an open-sourced, large-scale, and high-quality benchmark instruction dataset tailored for developing and evaluating LLMs in e-commerce realm. covers 116,528 samples from 10 real and widely performed e-commerce tasks of 4 categories. Each data sample comprises an instruction, an input, and an output. Some samples also incorporate a list of options. All the 10 tasks have in-domain (IND) test samples, and 6 tasks also have OOD test samples which consist of products unseen in the training samples of the respective tasks. undergoes rigorous and thorough scrutiny and is carefully crafted to enable a wide spectrum of empirical testing and exploration, including IND evaluation, OOD evaluation, and task-specific studies.
[…]
(1) models substantially outperform baseline models, including the most advanced GPT-4Β (OpenAI, 2023) and the state-of-the-art (SoTA) task-specific models, on almost all the 10 tasks in IND evaluation. On average, models show a substantial improvement of 10.7% over the best baseline models.
(2) Moreover, exhibits excellent generalizability to OOD settings, including unseen products and unseen instructions, highlighting its superiority as a generalist e-commerce model. Particularly, models establish an improvement of 9.3% over the best baselines on the OOD (new) products. These results indicate the great potential of both the dataset and the models in empowering versatile and effective LLMs for e-commerce, and validate the potential of LLMs in doing e-commerce tasks.
FigureΒ 1 depicts the overall scheme of . To the best of our knowledge, this study is the first comprehensive and systematic study of instruction-tuning LLMs for e-commerce applications, open-sources the first large-scale, high-quality benchmark dataset (), and develops the state-of-the-art generalist LLMs ( series) for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM/.
2 Related Work
Instruction Tuning with LLMsΒ Β Instruction tuning enables the transfer of general knowledge captured in LLMs to specific application domains, facilitating the generalizability of LLMs. FLANΒ (Wei etΒ al., 2021) and T0Β (Sanh etΒ al., 2021) are the early work to investigate instruction tuning by evaluating the zero-shot performance of fine-tuned LLMs on numerous datasets. Both of the models show encouraging performance over the original GPT-3Β (Brown etΒ al., 2020), indicating the effectiveness of instruction tuning. The importance of datasets, model scale, and instructions for instruction tuning are explored in a later work FLAN-v2Β (Chung etΒ al., 2024). Considering the significant impact of datasets for instruction tuning, several studies collect sizable benchmarks consisting of numerous tasks, such as Super-NaturalInstructionsΒ (Wang etΒ al., 2022b), Self-instructΒ (Wang etΒ al., 2022a) and Flan CollectionΒ (Longpre etΒ al., 2023), for fine-tuning general-purpose LLMs on NLP tasks. Meanwhile, instruction-tuned LLMs are employed in specific domains. For example, PlatypusΒ (Lee etΒ al., 2023a) is fine-tuned for reasoning in STEM, and MammothΒ (Yue etΒ al., 2024) is for general math problem-solving. Our incorporates instruction tuning with LLMs for e-commerce.
LLMs for E-commerceΒ Β LLMs are emerging in the general e-commerce realm. RecMindΒ (Wang etΒ al., 2023) utilizes LLMs as an autonomous recommender agent by enhancing their capability to comprehend user behaviors. LLaMA-EΒ (Shi etΒ al., 2023) is fine-tuned on LLaMAΒ (Touvron etΒ al., 2023a) for various e-commerce authoring tasks, such as ads generation and general product QA. EcomGPTΒ (Li etΒ al., 2024) is a pioneer instruction-tuned LLM on its dataset EcomInstruct. However, most of its tasks are human-constructed by repurposing data from other tasks (e.g., from the βquery product matchingβ task, a task is constructed to generate user queries based on the matched products), and thus, could have limited utility in real applications. A recent studyΒ (Geng etΒ al., 2022) develops a fine-tuned LLM, P5, with personalized prompts to conduct a few specific recommendation tasks (e.g., sequential recommendation). These studies demonstrate the utility of LLMs and instruction tuning in e-commerce applications. To the best of our knowledge, is the most comprehensive instruction-tuned e-commerce LLM.
Comparison among Instruction-tuned LLMs for E-commerceΒ Β TableΒ 1 summarizes the difference between our from LLaMA-E and EcomGPT, the two existing fine-tuned LLMs capable of performing multiple e-commerce tasks (P5 only performs recommendations). Fundamentally different from LLaMA-E and EcomGPT, focuses on comprehensive real-world e-commerce tasks and data, includes very extensive evaluation and benchmarking, and open-sources instruction data and models.
| Comparison | LLaMA-E | EcomGPT | |
| Β (Shi etΒ al., 2023) | Β (Li etΒ al., 2024) | (ours) | |
| Instruction tuning | β | β | β |
| Open-sourced data | β | β | β |
| Real-world tasks | β | β | β |
| Β Β Β Β Β # training tasks | 5 | 122β‘ | 10 |
| Β Β Β Β Β # In-domain tests | 5 | 0 | 10 |
| Β Β Β Β Β # out-of-domain tests | 0 | 12β‘ | 6 |
| # general-purpose | 5 | 3 | 5 |
| LLMs evaluated | |||
| # task-specific SoTA models evaluated | 0 | 0 | 11 |
| # base LLMs tuned | 3 | 4 | 6 |
| Open-sourced models | β | β | β |
β’
β‘EcomGPT has 122 training tasks, most of which are manipulated from data of different other tasks. The training data is not publicly available. EcomGPT releases 12 test tasks (8 in Chinese) for only out-of-domain evaluation.
3 Dataset
We introduce , an instruction dataset for adapting LLMs to e-commerce tasks. features 3 key design principles. (1) Broad coverage: includes 10 diverse tasks of 4 categories. Comprehensive and wide-ranging related tasks are critical in enabling versatile LLMs for a specific domain as shown in the literatureΒ (Lee etΒ al., 2023b). (2) Realistic tasks: focuses on real-world e-commerce tasks with real-world data, not human-manipulated tasks with synthetic data. This ensures e-commerce LLMs tuned on a potentially high utility in real-world applications. (3) High quality: considers only real-world data and undergoes rigorous scrutiny to ensure its accuracy and high quality. High data quality plays a pivotal role in building effective LLMsΒ (Hoffmann etΒ al., 2022; Gadre etΒ al., 2024).
3.1 E-commerce Tasks
comprises 10 real-world tasks with real-world data that are widely performed in e-commerce applications. These tasks fall within 4 categories: (i) product understanding, (ii) user understanding, (iii) query product matching, and (iv) product question answering. Such tasks are ubiquitous and essential on e-commerce platforms; success on these tasks would enable key functionalities in providing excellent user experience, promoting online retail, and driving sustainable revenues, among others.
Particularly, includes 3 tasks for product understanding: (1) attribute value extraction ()Β (Xu etΒ al., 2019; Wang etΒ al., 2020; Yang etΒ al., 2022), (2) product matching ()Β (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and (3) product relation prediction ()Β (Ahmed etΒ al., 2021; Xu etΒ al., 2020). For user understanding, includes (4) sentiment analysis ()Β (Wankhade etΒ al., 2022) and (5) sequential recommendation ()Β (Petrov & Macdonald, 2023; Li etΒ al., 2023). also covers 3 query product matching tasksΒ (Reddy etΒ al., 2022): (6) multi-class product classification (), (7) product substitute identification (), and (8) query-product ranking (). For product question answering, contains the tasks of (9) answerability prediction ()Β (Gupta etΒ al., 2019) and (10) answer generation ()Β (Deng etΒ al., 2023). More details are available in AppendixΒ A. TableΒ A1 summarizes the tasks and their data sources. Fundamentally different from the instruction data of LLaMA-EΒ (Shi etΒ al., 2023) and EcomGPTΒ (Li etΒ al., 2024), contains tasks all from real e-commerce platforms, and data all extracted from these real tasks.
3.2 Diverse Instructions
Recent workΒ (Xu etΒ al., 2022) shows that diverse instructions for LLM tuning improve LLM generalizability to new instructions. Inspired by this, we construct diverse instructions in . For each task, we start with a clear and concise human-written seed instruction. We then generate diverse instructions synonymous with the seed instruction via GPT-4Β (OpenAI, 2023), and select into 5 of those with writing styles (e.g., wording) distinct from the seed instruction. Thus, involves 6 high-quality and diverse instructions, including the seed instruction, on each task.
To test instruction understanding and model generalizability to new instructions, we hold out one instruction for each task as its βunseenβ instruction, which is inaccessible during model training. Thus, each task has 5 diverse instructions for instruction tuning, and 1 unseen instruction for testing. All the instructions are presented in AppendixΒ C.
3.3 Data Quality Control
In , we carry out the following procedures to ensure its accuracy and high quality. Specifically, we (1) remove overlapping data between training and test sets to avoid data leakage; (2) retain only data in English to ensure the unity of languages in texts; (3) eliminate non-English notations such as HTML tags and Unicode; (4) only select products with detailed information to allow sufficient product knowledge that LLMs can learn from; (5) keep texts within a reasonable length following the convention in the literatureΒ (Hou etΒ al., 2022); (6) manually inspect all processed data. We also conduct task-specific quality control on individual tasks. All the quality control procedures are detailed in AppendixΒ A.
3.4 Training, Validation, and Test Sets
is split into training sets, validation sets, IND test sets, and OOD test sets as follows. TableΒ A3 summarizes the dataset. AppendixΒ A describes the details of the data split process.
Training SetΒ Β has a training set of 10K for each individual task (except for , which has 2,022 training samples). The training sets from all the tasks are combined into a dataset of 92,022 samples as the training set of .
Validation SetΒ Β has a validation set of 1K for each individual task (except for , which has 253 validation samples). Similarly, the validation sets of each task are combined into a dataset of 9,253 samples as the validation set of .
In-Domain Test SetΒ Β For each of the 10 tasks, also includes an in-domain (IND) test set of 1K samples (except for , which has 253 IND samples). We define IND in terms of products that are within the same category as the training products.
Out-of-Domain Test SetΒ Β To evaluate the generalizability of LLMs tuned on to unseen samples β a critical capacity desired to address the cold-start problem in e-commerce, we create OOD test sets in . We define OOD in terms of products, that is, new products unseen during LLM training are considered as OOD. We determine OOD products using their category information. In , 6 tasks β , , , , , and , have different product categories. For each of these tasks, all products from one particular category are held out as the OOD data for that task, as summarized in TableΒ A3 (each product belongs to only one category). We focus on OOD (new) products, because promoting and recommending new products to users are effective strategies in e-commerce in driving sales and revenue, enhancing user experience, and increasing user engagement. Note we do not define OOD test sets over users (i.e., hold out users into OOD test sets), because users are anonymous and user identifiers are absent in existing datasets. Even though, the unseen instructions could be used in proximity to new users β a new user may have a completely new style of instructions to .
4 Models
We build a series of models on top of 6 base models: (1) : trained from large base models, Flan-T5 XXL (11B parameters)Β (Chung etΒ al., 2024) and Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (2) : trained from medium-sized base models, Llama-2 7B-chatΒ (Touvron etΒ al., 2023b) and Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023), and (3) : trained from small base models, Flan-T5 XL (3B)Β (Chung etΒ al., 2024) and Phi-2 (3B)Β (Javaheripi & Bubeck, 2023). These models are instruction-tuned over training data.
We fine-tune all the base models with LoRAΒ (Hu etΒ al., 2021) and Huggingface transformers libraryΒ (Wolf etΒ al., 2019). During fine-tuning, the learning rate and batch size of all the models are set as 1e-4 and 128, respectively. We apply a cosine learning rate scheduler with a 5% warm-up period for 3 epochs. We set and the rank in LoRA as 16, and add LoRA adaptors to all the projection layers and the language modeling head. We perform 0-shot evaluations (i.e., without in-context examples) on all the tasks.
5 Experimental Setup
| General-purpose LLMs | (1) GPT-4 TurboΒ (OpenAI, 2023), (2) Gemini ProΒ (Team etΒ al., 2023), (3) Claude 2.1Β (Anthropic, 2023), (4) Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (5) Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023) | |
| E-commerce LLM | (1) EcomGPTΒ (Li etΒ al., 2024): To the best of our knowledge, it is the only open-source e-commerce LLM. | |
| SoTA Task-specific Models | ||
| (SectionΒ A.1.1) | (1) SUOpenTagΒ (Xu etΒ al., 2019), (2)AVEQAΒ (Wang etΒ al., 2020): Both methods scan the product information, token by token, to extract the values of the specified attributes. | |
| (SectionΒ A.1.2) | (1) RGCNΒ (Schlichtkrull etΒ al., 2018): It utilizes a graph convolutional network to capture relations between products. (2) DeBERTaV3Β (He etΒ al., 2022): It predicts product relations from product titles. | |
| (SectionΒ A.2.1) | (1) BERTweetΒ (Nguyen etΒ al., 2020): It is a pre-trained language model for English tweets, and achieves superior performance over RoBERTaΒ (Liu etΒ al., 2019) and XLM-RΒ (Conneau etΒ al., 2019) on . (2) P5Β (Geng etΒ al., 2022): It is a pre-trained LLM for from e-commerce user reviews. | |
| (SectionΒ A.2.2) | (1) gSASRecΒ (Petrov & Macdonald, 2023), (2) RecformerΒ (Li etΒ al., 2023): Both methods leverage TransformerΒ (Vaswani etΒ al., 2017) to predict the next product of usersβ interest based on usersβ historical activities on products. | |
| (SectionΒ A.4.2) | (1) GPT-4 TurboΒ (OpenAI, 2023): To the best of our knowledge, there are no models specifically designed for . Thus, GPT-4 Turbo is used as the SoTA task-specific model in this task as it has outstanding performance in general question answeringΒ (OpenAI, 2023). | |
| , , , , (SectionΒ A.1.3, Β A.3.1, Β A.3.2, Β A.3.3, Β A.4.1). |
(1) BERTΒ (Devlin etΒ al., 2019), (2) DeBERTaV3Β (He etΒ al., 2022): Both methods generate predictions
from the textual product information (e.g., titles). |
We compare against 3 categories of baseline models: (1) general-purpose LLMs, (2) e-commerce LLMs, and (3) SoTA task-specific models. TableΒ 2 lists all the baseline models. Note that for the task-specific models, we only use the best (i.e., SoTA) models for each individual task based on literature, so as to enable schematic comparison between generalist modeling from and the task-specific modeling. We conduct IND and OOD tests on respective test datasets (SectionΒ 3.4) for all the models. TableΒ A3 lists the evaluations conducted for different tasks. Details on using/training baseline models are in AppendixΒ D. The prompt templates used in the evaluations are in AppendixΒ E.
General-purpose LLMsΒ Β For all the general-purpose LLMs, we use the checkpoints released by their developers. For GPT-4 Turbo, Gemini Pro, and Claude 2.1, we access the checkpoints using their official APIs. For Llama-2 13B-chat and Mistral-7B Instruct-v0.2, we use the checkpoints released in HuggingfaceΒ (Wolf etΒ al., 2019).
We perform 1-shot evaluations on all the general-purpose LLMs. Existing workΒ (Li etΒ al., 2024) tests LLMs on e-commerce tasks in a 0-shot setting. Meanwhile, in-context examples can notably benefit LLMsΒ (Brown etΒ al., 2020). Thus, in our case, we conduct 1-shot evaluations, balancing the computing cost and model performance, to enable stronger performance from the general-purpose LLMs. Although extensive prompt engineering and many in-context examples could enable better performance from LLMs, it is less practical in real e-commerce applications; for example, asking users to provide many in-context examples of the e-commerce platform, if ever feasible, can reduce user engagement. Also, it is not cost- and energy-efficient to do large-scale, few-shot in-context learning on LLMs over large test sets like βs.
E-commerce LLMsΒ Β For EcomGPT, we use the checkpoint released by its authors. We conduct both 0-shot and 1-shot evaluations on EcomGPT, as we empirically observe that 1-shot may result in better performance than 0-shot for EcomGPT (EcomGPT originally conducts 0-shot evaluations). Therefore, we report the best performance of the two evaluations on each task.
SoTA Task-specific Models We train the SoTA task-specific models SUOpenTag, AVEQA, RGCN, and gSASRec from scratch using training data of individual tasks and the model implementations published by their respective authors. For P5, we use the checkpoint released by its authors (it has already been pre-trained on e-commerce data). For GPT-4 Turbo, we access the checkpoint via its API. For Recformer, BERT, BERTweet, and DeBERTaV3, we tune the checkpoints on specific tasks.
| Model | ||||||||||
| F1* | Macro F1 | F1 | Macro F1 | HR@1 | Accuracy | F1 | NDCG | F1 | F | |
| GPT-4 Turbo | 0.495 | 0.326 | 0.753 | 0.516 | 0.387 | 0.611 | 0.195 | 0.875 | 0.649 | 0.858 |
| Gemini Pro | 0.396 | 0.136 | 0.867 | 0.470 | 0.269 | 0.584 | 0.248 | 0.821 | 0.506 | 0.855 |
| Claude 2.1 | 0.381 | 0.275 | 0.523 | 0.415 | 0.066 | 0.655 | 0.273 | 0.821 | 0.280 | 0.841 |
| Llama-2 13B-chat | 0.002 | 0.333 | 0.434 | 0.188 | 0.056 | 0.504 | 0.252 | 0.815 | 0.623 | 0.811 |
| Mistral-7B Instruct-v0.2 | 0.369 | 0.324 | 0.613 | 0.470 | 0.164 | 0.529 | 0.305 | 0.842 | 0.588 | 0.853 |
| EcomGPT | 0.000 | 0.091 | 0.648 | 0.188 | 0.042 | 0.540 | 0.170 | 0.000 | 0.086 | 0.669 |
| SoTA task-specific model | 0.546 | 0.588 | 0.995 | 0.573 | 0.265 | 0.703 | 0.389 | 0.859 | 0.830 | 0.858 |
| 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 | |
| 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 | |
| 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 | |
| improvement (%, avg: 10.7) | 21.2 | 3.9 | 0.0 | 13.1 | 40.1 | -1.0 | 28.8 | 0.1 | 2.5 | -1.9 |
β’
In this table, βF1*β, βMacro F1β, βF1β, βHR@1β, βAccuracyβ, βNDCGβ and βFβ are the primary evaluation metrics in respective tasks (AppendixΒ A). For each task, the best baseline performance is underlined, and the overall best performance is in bold. The row βimprovementβ presents the percentage improvement of the best-performing model over the best-performing baseline model (underlined) in each task. We also include the average (βavgβ) improvement across all the tasks in the table.
6 Experimental Results
We evaluate the models using the test sets of individual tasks. For each task, we employ multiple metrics for a comprehensive evaluation (AppendixΒ A). For the sake of clarity, in this section, we present the performance only in terms of the primary evaluation metric for each task (TableΒ A1). As discussed in AppendixΒ F, Llama-2 13B-chat, Mistral-7B Instruct-v0.2, and Phi-2 are the best base models for , , and , respectively. Thus, by default, , , and refer to those tuned from these base models, respectively.
Main ResultsΒ Β Our comprehensive experiments yield the following main results: (1) models demonstrate the best performance on almost all the IND tasks, with a significant average improvement of 10.7% over the general-purpose LLMs, e-commerce LLMs, and the SoTA task-specific models across the 10 tasks (SectionΒ 6.1). (2) models show outstanding generalizability to OOD products and surpass the best baselines with a remarkable average improvement of 9.3% in OOD evaluation (SectionΒ 6.2). (3) By training over diverse instructions, is equipped with strong generalizability to unseen instructions (SectionΒ 6.3). (4) Trained on all the tasks in together, exhibits similar or better performance than models trained on each individual task (SectionΒ 6.4). (5) models benefit from larger instruction training data for e-commerce tasks (SectionΒ 6.5).
6.1 In-domain Evaluation
TableΒ 3 shows the performance of and all the baseline methods in IND evaluation, where models are trained using training set (i.e., including all the tasks), and the SoTA task-specific models are trained using the task-specific training data. Among the SoTA task-specific models, we report the results of only the best-performing model on each task. The complete results of each task are presented in AppendixΒ G.
Overall ComparisonΒ Β As shown in TableΒ 3, overall, substantially outperforms baseline models across the 10 e-commerce tasks at 10.7% on average. Particularly, models (i.e., , , and ) achieve superior performance over the baselines on 7 out of the 10 tasks with an average improvement of 15.7%. On the rest 3 tasks , , and , models achieve the same or comparable performance as the baselines (e.g., maximum difference of 1.9% as on ). These results demonstrate the remarkable effectiveness of compared with the general-purpose LLMs, the SoTA task-specific models, and the existing e-commerce LLM across the e-commerce tasks.
Comparison between and General-purpose LLMsΒ Β TableΒ 3 shows that models substantially outperform the general-purpose LLMs by a remarkable margin. For example, across the 10 tasks, achieves a significant average improvement of 39.6% over GPT-4 Turbo. A key difference between and GPT-4 Turbo is that is specifically tuned on our instruction dataset for e-commerce. The remarkable improvement of over general-purpose LLMs suggests that there could be a significant gap between general knowledge and the knowledge required for e-commerce tasks, highlighting the significance of in imparting knowledge pertinent to e-commerce into LLMs.
We also observe that general-purpose LLMs lag behind the SoTA task-specific models by a large margin. For example, the best general-purpose LLM GPT-4 Turbo considerably underperforms the SoTA task-specific models on 7 out of the 10 tasks (e.g., 0.495 vs 0.546 on ). This underscores the critical need to deliberately accommodate general-purpose LLMs for e-commerce tasks.
Comparison between and E-commerce LLMsΒ Β According to TableΒ 3, models demonstrate substantial improvement over the existing e-commerce LLM EcomGPT on all the tasks. For example, surpasses EcomGPT remarkably by 244.6% on , and by 53.5%. Both and EcomGPT are instruction-tuned LLMs for e-commerce. However, our instruction dataset for fundamentally differs from EcomGPTβs EcomInstruct dataset: includes only real-world data and real-world e-commerce tasks, while EcomInstruct incorporates a considerable amount of synthetic data and tasks, which hinders its applicability in real e-commerce tasks.
Comparison between and SoTA task-specific modelsΒ Β TableΒ 3 shows the SoTA task-specific models perform best among the baselines on each respective task. The substantial improvement of over the SoTA task-specific models serves as strong evidence that , with specific tuning over , could leverage knowledge shared across multiple e-commerce tasks and boost performance on each individual task.
6.2 Out-of-domain Evaluation
| Model | ||||||
| F1* | M-F1 | M-F1 | HR@1 | F1 | F | |
| GPT-4 Turbo | 0.397 | 0.392 | 0.510 | 0.198 | 0.680 | 0.860 |
| Gemini Pro | 0.275 | 0.123 | 0.454 | 0.116 | 0.552 | 0.856 |
| Claude 2.1 | 0.410 | 0.277 | 0.369 | 0.036 | 0.245 | 0.842 |
| Llama-2 13B-chat | 0.000 | 0.324 | 0.178 | 0.050 | 0.644 | 0.808 |
| Mistral-7B | 0.264 | 0.327 | 0.438 | 0.108 | 0.608 | 0.851 |
| Instruct-v0.2 | ||||||
| EcomGPT | 0.001 | 0.096 | 0.178 | 0.023 | 0.140 | 0.722 |
| SoTA | 0.269 | 0.507 | 0.567 | 0.081 | 0.853 | 0.860 |
| task-specific model | ||||||
| 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 | |
| 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 | |
| 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 | |
| improvement | -10.5 | 10.1 | 14.1 | 41.4 | 3.0 | -2.2 |
| (%, avg: 9.3) |
β’
In this table, βM-F1β represents macro F1. The columns in this table have the same meanings as those in TableΒ 3.
TableΒ 4 shows the performance of and baselines in OOD evaluation. Overall, we observe a similar trend to that in IND evaluation. Specifically, outperforms the SoTA task-specific models on OOD evaluation by a wide margin on 5 out of the 6 tasks, except for task, on which is comparable with the SoTA task-specific model (e.g., 2.2% difference). Similarly, outperforms the general-purpose LLMs on 4 out of the 6 tasks (i.e., , , , and ). Across the 6 tasks, demonstrates a substantial average improvement of 18.9% over GPT-4 Turbo. Compared to EcomGPT, again shows terrific advantages on all the tasks in OOD evaluation. As shown in the literatureΒ (Lika etΒ al., 2014), cold start for new products has been an unsolved issue but a key driver in e-commerce applications. The OOD generalizability of to new products as demonstrated in TableΒ 4 makes a highly viable tool for e-commerce applications.
Note that GPT-4 and Claude 2.1 exhibit robust performance on and , indicating that these general-purpose LLMs are inherent with rich knowledge for solving extraction and question-answering problems. However, they may still lack comprehensive knowledge to effectively perform multiple, diverse e-commerce tasks, such as on the other 4 tasks.
6.3 Generalizability to Unseen Instructions
| Model | Training Instructions | ||||||||||
| F1* | Macro F1 | F1 | Macro F1 | HR@1 | Accuracy | F1 | NDCG | F1 | F | ||
| single | 0.046 | 0.619 | 0.995 | 0.610 | 0.526 | 0.696 | 0.206 | 0.870 | 0.846 | 0.841 | |
| diverse | 0.553 | 0.638 | 0.995 | 0.639 | 0.524 | 0.694 | 0.335 | 0.870 | 0.842 | 0.841 | |
| single | 0.000 | 0.618 | 0.995 | 0.554 | 0.543 | 0.696 | 0.241 | 0.878 | 0.852 | 0.850 | |
| diverse | 0.622 | 0.540 | 0.995 | 0.643 | 0.540 | 0.695 | 0.253 | 0.878 | 0.822 | 0.844 | |
| single | 0.447 | 0.535 | 0.991 | 0.577 | 0.478 | 0.652 | 0.314 | 0.867 | 0.841 | 0.838 | |
| diverse | 0.488 | 0.552 | 0.991 | 0.577 | 0.457 | 0.660 | 0.381 | 0.871 | 0.845 | 0.842 |
β’
In this table, βsingleβ and βdiverseβ indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model (e.g., , abd ) is in bold, if the performance difference between the model tuned over single and diverse instructions exceeds 1%.
| Model | Training Tasks | ||||||||||
| F1* | Macro F1 | F1 | Macro F1 | HR@1 | Accuracy | F1 | NDCG | F1 | F | ||
| Task-specific | 0.599 | 0.521 | 0.995 | 0.616 | 0.518 | 0.655 | 0.000 | 0.879 | 0.854 | 0.841 | |
| Generalist | 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 | |
| Task-specific | 0.757 | 0.543 | 0.987 | 0.655 | 0.535 | 0.681 | 0.000 | 0.883 | 0.864 | 0.841 | |
| Generalist | 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 | |
| Task-specific | 0.397 | 0.348 | 0.991 | 0.608 | 0.413 | 0.646 | 0.000 | 0.858 | 0.835 | 0.835 | |
| Generalist | 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 |
β’
In this table, βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold, if the performance difference between the generalist and task-specific model exceeds 1%.
To evaluate the generalizability of to unseen instructions, we compare models tuned from diverse instructions and tuned from single instructions per task, and test their performance over unseen instructions.
As shown in TableΒ 5, overall, models tuned over multiple, diverse instructions exceed or resemble those tuned over single instructions. For example, performs much better on unseen instructions on and when trained from diverse instructions. On the other 8 tasks, trained from diverse instructions is either better than or comparable to that from single instructions. Similarly, in the case of , using diverse instructions shows very strong effects on performance on and , while for other tasks, performance remains comparable (except for ). For , diverse instructions enhance performance on almost all the tasks.
These results illustrate the pivotal role of diverse instructions for e-commerce tasks, and underscore the utility of in generalizing LLMs to new instructions in e-commerce applications. The results also showcase the generalizability of to unseen instructions. As new users may apply instructions unseen in existing e-commerce data, such generalizability also indicates the potential of for cold-start users in e-commerce. It is noticeable that is able to generalize to unseen instructions (TableΒ 5) with a similar performance as it has on βseenβ instructions (TableΒ 3). For example, on , shows only a 1.4% difference in its performance in the former setting and in the latter setting (0.639 vs 0.648). This further demonstrates the strong generalizability of .
6.4 Generalist vs Task-specific Models
We compare the models tuned using all tasks (generalist models) with those tuned on individual tasks (task-specific models) and present the performance comparison in IND evaluation in TableΒ 6. The results of the OOD evaluation are presented in AppendixΒ H.3. As shown in TableΒ 6, generalist demonstrates slightly stronger performance compared to task-specific on each task (e.g., 2.7% average improvement of generalist over all tasks except ). For example, on , generalist is comparable to the task-specific (0.582 vs 0.599). On , generalist demonstrates a substantial improvement of 17.3% compared to task-specific (0.611 vs 0.521). As evidenced by the results, by training on all tasks together, models enjoy strong versatility and could enable knowledge transfer across tasks for improved performance.
6.5 Impact of Training Data Size on
We investigate how the training data size impacts the performance of models. As presented in SectionΒ 3.4, has more than 92 thousand (92K) training samples. In this experiment, we generate 3 smaller training sets of 1K, 10K, and 47K samples. The 1K and 10K samples are generated by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are constructed by randomly selecting 5K samples from each task except for , for which we include all its 2K training samples. The complete results are presented in AppendixΒ I.
FigureΒ 2 presents the model performance on . As it shows, increasing training data size typically benefits model performance. For example, in IND evaluation, the performance of increases from 0.085 to 0.526 when the training data size is increased from 1K to 92K. Similarly, in OOD evaluation, increasing training samples from 1K to 92K significantly elevates performance from 0.059 to 0.273. A similar trend could be observed in and . As evidenced by our results, the large-scale training data is critical in developing effective e-commerce LLMs. This further highlights the significance of our extensive, comprehensive, and high-quality e-commerce instruction dataset, .
7 Conclusion and Limitation
This paper open-sources the first large-scale, high-quality benchmark dataset (), and develops the state-of-the-art generalist LLMs ( series) for e-commerce. models are extensively evaluated against the most advanced baseline models including GPT-4 Turbo, EcomGPT, and SoTA task-specific models, in both IND and OOD evaluations. Our experimental results demonstrate that substantially outperforms the best baseline models with an average improvement of 10.7% in IND evaluation. Our results also show that exhibits excellent generalizability to OOD settings, evidenced by its considerable improvement of 9.3% over the best baselines on the OOD products. To the best of our knowledge, this study is the first comprehensive and systematic study of instruction-tuning LLMs for e-commerce applications.
This paper acknowledges certain limitations and future work. (1) Recently emerged e-commerce tasks such as explanation generation could be included to further improve the comprehensiveness of . (2) User profiling could be better enabled once metadata is available, and advanced foundation models could be developed to address unique and fundamental challenges and tasks in e-commerce.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Task Definition and Data Preprocessing
All tasks can be characterized into 4 categories: product understanding, user understanding, query product matching, and product question answering (product QA). TableΒ A1 summarizes the task definitions, evaluation metrics, data sources for each task, and their OOD test settings, if any. The details are articulated below.
| Task | Definition | Type | Metrics | Data | |
| Product Understanding | Given the titles, descriptions, features, and brands of the products, extract values for the specific target attributes. | Information extraction | precisionβ, recallβ, F1β (SectionΒ A.1.1) | MAVEΒ (Yang etΒ al., 2022) based on Amazon Review 2018Β (Ni etΒ al., 2019); OOD: 7 held-out attributes | |
| Given the titles of two products, predict their relation from βalso buyβ, βalso viewβ, and βsimilarβ. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category | ||
| Given the titles, descriptions, manufacturers, and prices of the products from two different platforms, predict if they are the same product. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | Amazon-Google ProductΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) | ||
| User Understanding | Given a product review by a user, identify the sentiment that the user expressed on the product. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category | |
| Given the interactions of a user over the products, predict the next product that the user would be interested in. | Ranking | HR@1 | Amazon Review 2018Β (Ni etΒ al., 2019) and Amazon Review 2014Β (He & McAuley, 2016; McAuley etΒ al., 2015); OOD: Tools category | ||
| Query Product Matching | Given a query and a product title, predict the relevance between the query and the product. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | |
| Given a user query and a potentially relevant product, predict if the product can serve as a substitute for the userβs query. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | ||
| Given a user query and a list of potentially relevant products to the query, rank the products according to their relevance to the query. | Ranking | NDCG[1] | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | ||
| Product QA | Given a product-related question and reviews of this product, predict if the question is answerable. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category | |
| Given a product-related question and reviews as supporting documents, generate the answer to the question. | Generation | P[2], R[2], F[2], BLEURT[3] | AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category |
β’
Underline indicates the primary metrics. Higher values of the primary metrics indicate better model performance. OOD refers to the out-of-domain product category/attributes. [1] NDCG: normalized discounted cumulative gainΒ (Wang etΒ al., 2013) [2] P, R, F: BERTScore evaluates the quality of generated texts by measuring the similarity between the embeddings of tokens in the generated text and the ground-truth textΒ (Zhang etΒ al., 2019). [3] BLEURT measures the text quality by comparing the similarity of sentences using contextual embeddings from pre-trained models. Β (Sellam etΒ al., 2020).
To pursue adherence to data usage requirements, we check the licenses of data sources, ensuring their permission to publish. Table A2 presents the licenses of our curated dataset sources.
| Dataset | License Type | Source |
| Amazon-Google Products | CC-by-4.0 | https://dbs.uni-leipzig.de/research/projects/benchmark-datasets-for-entity-resolution |
| Amazon Review | Not Specified | https://cseweb.ucsd.edu/jΜmcauley/datasets.html#amazon_reviews |
| AmazonQA | Not Specified | https://github.com/amazonqa/amazonqa |
| Shopping Queries Dataset | Apache License 2.0 | https://github.com/amazon-science/esci-data |
Data Split
Raw datasets of the attribute value extraction (, discussed below in Section A.1.1), product matching (, discussed in Section A.1.3), product relation prediction (, discussed in Section A.1.2), and sentiment analysis (, discussed in Section A.2.1) tasks are first split into training, validation, and test data at 8:1:1 ratio. For multi-class product classification (, discussed in Section A.3.1), query product ranking (, discussed in Section A.3.3), product substitute identification (, discussed in Section A.3.2), answerability prediction (, discussed in Section A.4.1), and answer generation (, discussed in Section A.4.2) tasks, the raw datasets are already split. All split data are processed as detailed below and converged with instruction templates into structured data, which will be directly used by LLMs. Based on prior researchΒ (Wei etΒ al., 2021) and considering the high computing demands, we uniformly downsample training sets to a size of 10K, validation sets to 1K, and test sets to 1K size, optimizing data volumes for efficient processing and affordable LLM evaluation.
In-domain (IND) Data Selection
In , , and tasks, we utilize the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) as the data source, which contains product reviews and metadata in 29 different product categories. For these three tasks, the data from the Electronics, Home, and Sports categories are employed as the IND data sources. We further downsample and combine the processed structured data for each task from these sources at a 1:1:1 ratio into a training, validation, and test set, respectively. For and tasks, we use the data from the AmazonQA datasetΒ (Gupta etΒ al., 2019), which has 17 product categories. To avoid data leakage, we use the data from the Sports and Tools categories for the task, and Electronics and Home for task as IND data sources, respectively. Similarly to the previous three tasks, for each of the tasks, we sample and combine the processed data from the IND sources at a 1:1:1 ratio into training, validation, and test sets. For the task, we process all data from Amazon-Google ProductsΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and get a 2K training set, 0.2K validation set, and 0.2K test set. Since this dataset is small, we do not do the downsampling for the task. For , , , and tasks, we directly process and downsample them from the corresponding split in original datasets, separately.
Out-of-domain (OOD) Data Selection
As for OOD datasets, only test data are utilized. The data from the Tools category from the Amazon Review dataset is used as the OOD data source in , , and tasks, while the Cells category from AmazonQA serves as the OOD source in and tasks. For the task, we hold out 7 attributes as the OOD dataset. All OOD sources are processed into 1K test sets.
| Task | Training | Validation | IND Test | OOD Test |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 02,022 | 0,253 | 0,253 | β | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 92,022 | 9,253 | 10 tasks | 6 tasks |
β’
In this table, IND and OOD refers to the in-domain evaluation and out-of-domain evaluation, respectively.
A.1 Product Understanding Tasks
Three tasks are defined as follows to understand the different aspects of products.
A.1.1 Attribute Value Extraction ()
Definition:
Given the titles, descriptions, features, and brands of the products, extract values for the specific target attributes. By understanding product attribute values, models can extract key properties of the products and build the profiling for them, which is beneficial in many e-commerce scenarios, such as customer service agents and explanations.
Data Processing:
The Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) is used as the raw data. We extract the ground-truth attribute-value pairs following the protocol of MAVEΒ (Yang etΒ al., 2022) and get 4,767,579 data entries with 683 attributes in total. We randomly hold out 7 attributes from the dataset as the OOD data source and process both the OOD data and remaining IND data with instructions. The samples that have ground-truth values and corresponding sources for the specific attribute in the dataset are referred to as positive samples, while the samples that have no extracted ground-truth values are denoted as negative samples.
Evaluation Metrics:
Following the evaluation definition in previous work MAVEΒ (Yang etΒ al., 2022), the model can predict null value (NV) or incorrect value (IV) for negative samples. For positive samples, the model predictions can be correct value (CV), wrong value (WV), and null value (NL). In MAVE, customized precision and recall are calculated as the percentage of correctly predicted positive samples in all predictions and ground truths, respectively. This formulation, however, fails to evaluate model results on negative samples. Therefore, with the definition above, we improve the formulation by considering both positive and negative samples as follows:
where N refers to the total number of entries in ground-truth data. Note that we only consider the prediction result as the correct value (VC) when the ground truth is fully contained in the prediction (e.g. βbright yellowβ is considered as a correct value with the ground truth of βyellowβ). In our formulation, the precisionβ is calculated as the percentage of correctly predicted positives and negatives in all predictions, and the recallβ is computed as the percentage of correct positive and negative predictions in all ground truths. We use F1β as the primary metric for the AVE task.
Example:
β’
Input:
β
Product title: Bencore Multi Functional Molle Tactical Messenger Bag.
β
Product description: This rugged/durable tactical shoulder bag provides perfect and stylish solution for almost any scenario. The bag is made of durable nylon construction that will not tear, color will not fade. The bag has many MOLLE straps through the bag for all your MOLLE accessories. The bag contains many roomy compartments as pictured and comes in many stylish colors. Design, comfort and functionality was the emphasis of this bag which is why we made sure the bag is fully ergonomic, lightweight and has many roomy pockets and Velcro patches throughout the bag. The product comes with the Bencore Life Time warranty and is satisfaction guaranteed. Bencore is a leading manufacturer in outdoor Apparel/Accessories, from Par cords to Backpacks to basic outdoor essentials.
β
Product feature: Durable heavy-duty, lightweight Nylon construction, will not tear or break even under extreme conditions - Lifetime Warranty, Rugged, roomy main drawstring-closed compartment provides secure storage space for your gear; MOLLE System, works with most MOLLE accessory, Front pocket provides quick access, roomy interior pocket for convenient separated storage, concealed back pocket with zipper closure, Padded and fully ergonomic System, adjustable shoulder strap for comfortable handling.
β
Product brand: Bencore
β
Target attributes: Material
β
β’
Output:
β
Attribute: material; Value: nylon; Source: product description.
β
Attribute: material; Value: nylon; Source: product feature.
β
Note that in the above example, for the βtarget attributeβ βmaterialβ, the model should extract its value βnylonβ from the βproduct descriptionβ and βproduct featureβ.
A.1.2 Product Relation Prediction ()
Definition:
Given the titles of two products, predict their relation. Studying the relations between products can help models generate better results when conducting other e-commerce tasks such as recommendations.
Data Processing:
To learn the relationship between products, we use the product metadata of Electronics, Home, and Sports categories from the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) as the IND sources and Tools as the OOD source. We collect product IDs from metadata and remove the products without detailed information in the metadata. In this task, the product titles are used to represent products. The product pairs that appear more than once with different relations are eliminated. After filtering and combining data with instruction templates, the three relations (also buy, also view, and similar) in the structured dataset are roughly 7:10:1.
Evaluation Metrics:
With three different relations in this task, accuracy, macro precision, macro recall, and macro F1 are employed as metrics. Meanwhile, macro F1 is used as the primary metric for combining evaluation results of different labels and providing a comprehensive measurement of the model performances.
Example:
β’
Input:
β
Product 1: Monoprice 11952 Polyurethane Replacement Ear Pads for PID 8323 type Headphones - Red
β
Product 2: Monoprice Hi-Fi Light Weight Over the Ear Headphones - Black with a 50mm driver and a 47in 3.5mm cable for Apple iPhone iPod Android Smartphone Samsung Galaxy Tablets MP3
β
β’
Options:
β
A. Users who view product 1 may also buy product 2.
β
B. Users who view product 1 may also view product 2.
β
C. The product 1 is similar with the product 2.
β
β’
Output:
β
B
β
Note that in the above example, the relation of Product 1 and Product 2 is also_view.
A.1.3 Product Matching ()
Definition:
Given the titles, descriptions, manufacturers, and prices of the products from two different platforms, predict if they are the same product. This task enables the model to learn the similarities among products.
Data Processing:
We use the original dataΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010), which contains detailed product information from Amazon and Google platforms, as well as 1.3K matching product pairs between the two platforms. After deduplication, we randomly sample the same amount of unmatched pairs since there are only matched samples in the raw data and thus get 2,530 product pairs in total. The product pairs are split into the training, validation, and test sets with a ratio of 8:1:1, and processed with instruction templates respectively.
Evaluation Metrics:
For this binary classification task, we use precision (positive prediction rate), recall (sensitivity), F1, specificity (recall of negative labels), negative prediction rate (NPR, precision of negative labels), and accuracy as metrics. We choose F1 as the primary metric since it serves as a balanced evaluation metric by combining precision and recall to generally reflect model performance.
Example:
β’
Input:
β
Product 1: title - marine aquarium 2.5 virtual undersea paradise win/mac, description - marine aquarium 2.0 is like having a small piece of an aquatic paradise in your home -- without having to take care of actual fish, manufacturer - encore software, price - 19.99
β
Product 2: title - encore software 25020 - marine aquarium 2.5 (hybrid) - win 95 98 me 2000 xp/mac 10.1 or higher, description - encore software 25020: marine aquarium 2.5 hybrid discover the virtual fish tank phenomenon that has everyone talking! marine aquarium 2.5 delivers a stunning undersea paradise through your desktop with 26 exotic species of fish, manufacturer - encore software, price - 19.97
β
β’
Output:
β
Yes
β
In the above example, the two products (Product 1, Product 2) are identified as the same product.
A.2 User Understanding
The two tasks in this section aim to help the models comprehend the usersβ needs and preferences.
A.2.1 Sentiment Analysis ()
Definition:
Given a product review by a user, identify the sentiment that the user expressed on the product. The task will help models understand what sentiment users express and recommend more proper products to the user
Data Processing:
For the sentiment analysis, we also use the review data of Electronics, Home, and Sports categories from the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) as the IND sources and Tools category as the OOD source. We only keep the data with reviews of at least 10 words. The ratings are used as the ground-truth sentiment level, while 5.0 refers to very positive and 1.0 refers to very negative. After downsampling and combining the raw data with instruction templates, the five labels (from 1.0 to 5.0) in the structured dataset are roughly 4:2:3:8:25.
Evaluation Metrics:
With five different labels in the task, accuracy, macro precision, macro recall, and macro F1 are employed as metrics. macro F1 is used as the primary metric.
Example:
β’
Input:
β
This is really perfect for my kids who have a thick hair. I cam be able to create a beautiful hair bun with them. I would like to recommend this to all.
β
β’
Options:
β
A. Very positive
β
B. Positive
β
C. Neutral
β
D. Negative
β
E. Very negative
β
β’
Output:
β
A
β
The user from the above example expressed a very positive sentiment in the review.
A.2.2 Sequential Recommendation ()
Definition:
Given the interactions of a user over the products, predict the next product that the user would be interested in. By learning on this task, the models will have a comprehensive view of user preferences, which enables models to cater to usersβ future needs.
Data Processing:
In the task, we use both product reviews and metadata from the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019). Meanwhile, we use metadata from the Amazon Review 2014 datasetΒ (He & McAuley, 2016; McAuley etΒ al., 2015) as a supplement. The data of Electronics, Home, and Sports categories from both datasets serve as IND sources and the Tools category is used for the OOD test. Moreover, we consider usersβ review histories as their interactions with products. Following the data processing protocol of UnisRecΒ (Hou etΒ al., 2022), we remove the products without metadata in the 2018 version dataset and conduct the 5-core filter to ensure all products and users appear at least 5 times. After filtering, we sort every userβs interactions chronologically and truncate the history with a maximum of 50 products, retaining the least recent history. For the text information of products, we use the metadata from Amazon Review 2014 to fill in the missing information of the same products in the 2018 version metadata.
We also combine the product title, category, and brand to represent a product. The average length of the combined texts is about 21 words. Thus, we retain the first, maximum of 25 words of each combined text for computational efficiency. As in conventional sequential recommendation tasks, we split the last product of the user interactions into the test set as the ground truth next product of the userβs interest, the second last product into the validation set, and the remaining products into the training set. When processing the user interactions with instruction templates, for each sample, we randomly select 19 candidate products and mix them up with one ground-truth product. These 20 products serve as the options for the sample.
Evaluation Metrics:
We evaluate the task on hit rate at top 1 (HR@1), which is a popular metric in sequential recommendation and measures if the top-ranked product matches the ground-truth user interaction. HR@1 also serves as the primary metric in this task.
Example:
β’
Input:
β
1st: M-Edge Latitude Kindle Jacket, Pink (Fits Kindle Keyboard). Electronics. Computers & Accessories. M-Edge.
β
2nd: Marware jurni Kindle Fire Case Cover, Black (will not fit HD or HDX models). Electronics. Computers & Accessories. Marware.
β
3rd: NETGEAR AC1600 Dual Band Wi-Fi Gigabit Router (R6250). Electronics. Computers & Accessories. NETGEAR.
β
4th: iMBAPrice 110014-1 (1-Pack) Glod Plated 2.4 Ghz 3-Way Coaxial Cable Splitter F-Type Screw for Video Satellite Splitter/VCR/Cable Splitter/TV Splitter/Antenna Splitter/RG6 Splitter. Electronics. Accessories & Supplies...
β
β’
Options:
β
A: T POWER 9vΒ 12v (6.6ft Long Cable) Ac Dc Adapter Compatible with X Rocker Pro Series H3 51259 Video Gaming Chair 51231,51396 & V Rocker 5130301...
β
B: Boys Floatsafe Flotie Soft Fabric Armbands Floatie Blue For Kids Ages 1 To 3. Floatsafe Floatie
β
C: Anker iPhone Charger, Powerline Lightning Cable (3ft), MFi Certified for iPhone Xs/XS Max/XR/X
β
D: Curtain Drapery Rod w/brackets Small - Wrought Iron Hand Made. Home & Kitchen. Home Dcor. Hand Crafted & American Made!
β
β¦
β
T: Lorex ACCMIC1 Indoor Audio Microphone Accessory for Surveillance DVRβs (Black). Electronics. Camera & Photo. Lorex
β
β’
Output:
β
A
β
Given the interactions shown in the above input, the next product of userβs interest for this user is C: Anker iPhone Charger, Powerline Lightning Cable (3ft), MFi Certified for iPhone Xs/XS Max/XR/X
A.3 Query Product Matching
The following three tasks seek to study the relations between the user queries and the potential relevant products to the queries. We use the data from the Shopping Queries DatasetΒ (Reddy etΒ al., 2022), which contains 48,300 queries and potentially relevant products to each query. The dataset was originally collected from different market locales in various languages. All products are labeled in four categories: Exact, Substitute, Complement, and Irrelevant, based on their relevance to the query. As in the Shopping Queries Dataset, the four labels are defined as follows:
β’
Exact: The item is relevant for the query, and satisfies all the query specifications (e.g., a water bottle matching all attributes of a query βplastic water bottle 24ozβ, such as material and size).
β’
Substitute: The item is somewhat relevant, i.e., it fails to fulfill some aspects of the query, but the item can be used as a functional substitute (e.g., fleece for a βsweaterβ query)
β’
Complement: The item does not fulfill the query, but could be used in combination with an exact item (e.g., track pants for βrunning shoesβ query).
β’
Irrelevant: The item is irrelevant, or it fails to fulfill a central aspect of the query (e.g., socks for a βtelescopeβ query, or a wheat flour bread for a βgluten-free breadβ query)
In the following query-related tasks, we only utilize the data that is both in English and from the market of the U.S. locale.
A.3.1 Multi-class Product Classification ()
Definition:
Given a query and a product title, predict the relevance between the query and the product (Exact, Substitute, Complement, Irrelevant). This task helps models learn the fine-grained relevance between queries and products, promoting better recommendation results.
Data Processing:
We use the product titles to represent products and translate the Unicode into English. The translated query-product pairs are processed with instruction templates. The ratio of the four labels (Exact, Substitute, Complement, and Irrelevant) in the structured dataset is about 35:10:1:5.
Evaluation Metrics:
Accuracy, macro precision, macro recall, and macro F1 are employed as metrics. Accuracy is used as the primary metric in the task.
Example:
β’
Input:
β
Query: aj1 black and white
β
Product: Nike Menβs Air Jordan 1 Low White/Gym Red, White/Gym Red/Black, 9
β
β’
Options:
β
A: The product is relevant to the query, and satisfies all the query specifications.
β
B: The product is somewhat relevant. It fails to fulfill some aspects of the query but the product can be used as a functional substitute.
β
C: The product does not fulfill the query, but could be used in combination with a product exactly matching the query.
β
D: The product is irrelevant to the query.
β
β’
Output:
β
B
β
In the above example, the product serves as a substitute for the product described in the query.
A.3.2 Product Substitute Identification ()
Definition:
Given a user query and a potentially relevant product, predict if the product can serve as a substitute for the userβs query.
Data Processing:
The preprocessing of the is similar to that of , except that the PSI is a binary classification task. The query-product pairs with Exact, Complement, or Irrelevant labels are relabeled as non-substitute. After combining the query-product-label triples with instruction templates, the ratio of substitute (positive) and non-substitute (negative) labels is approximately 4:1.
Evaluation Metrics:
For this binary classification task, we use precision (positive prediction rate), recall (sensitivity), F1, specificity (recall of negative labels), negative prediction rate (NPR, precision of negative labels), and accuracy as metrics. We choose F1 as the primary metric.
Example:
β’
Input:
β
Query: fissler magic smooth-edge can opener
β
Product: KUKINO Manual Can Opener, Multifunction Handheld Food Grade Stainless Steel Can Openers, Black.
β
β’
Output:
β
No
β
The product does not serve as a substitute for the query in the above example.
A.3.3 Query-product Ranking ()
Definition:
Given a user query and a list of potentially relevant products to the query, rank the products according to their relevance to the query.
Data Processing:
In the task, we utilize the product titles to represent the products. A query and a list of potentially relevant products with different relevance labels to this query constitute a sample. We sort the list of products according to their relevance and generate the ground truth for each sample.
Evaluation Metrics:
We evaluate this task using normalized discounted cumulative gain (NDCG)Β (Wang etΒ al., 2013), which is a prevalent evaluation metric for assessing the ranking quality in the existing literature. NDCG is also used as the primary metric in this task.
Example:
β’
Input:
β
Query: high heel shoe chair
β
Product A: ORE International HBB1826 High Heel Shoe Display with Hooks Jewelry Box, Cheetah Print.
β
Product B: Coconut Float Red High Heel Gigantic Pool Float for Adults, 91.
β
Product C: Wildkin Kids Wooden Bench Seat with Storage for Boys and Girls, Toy Box Bench Seat Features Safety Hinge, Backrest, and Two Carrying Handles, Measures 32 x 15.5 x 27 Inches (Wild Side) (LOD71001).
β
β’
Output:
β
A, C, B
β
In the above example, Product A is more relevant to the query than Product C, and Product C is more relevant than Product B.
A.4 Product Question Answering
These two tasks aim to learn from the products and answer the product-related questions by using the AmazonQA datasetΒ (Gupta etΒ al., 2019). This dataset involves 923,685 product-related questions. There are around 9 product-related reviews and 4 answers provided by Amazon users for each question. There are also questionType and is_answerable annotations for each question. We eliminate all meaningless notations such as HTML tags or emoji from questions, reviews, and answers.
A.4.1 Answerability Prediction ()
Definition:
Given a product-related question and reviews of this product, predict if the question is answerable.
Data Processing:
We use the data of the Sports and Tools categories from AmazonQA as the IND sources and the Cell category as the OOD source. The is_answerable annotations are used as the ground truth. The ratio of answerable (positive) and unanswerable (negative) samples in the structured dataset is around 2:3.
Evaluation Metrics:
For the task, we use precision (positive prediction rate), recall (sensitivity), F1, specificity (recall of negative labels), negative prediction rate (NPR, precision of negative labels), and accuracy as metrics. We use F1 as the primary metric when evaluating this task.
Example:
β’
Input:
β
Question: Where do you purchase the paddles or do paddles come with it?
β
[…]
β
β’
Output:
β
No
β
From the above example, the question is unanswerable based on the document.
A.4.2 Answer Generation (AG)
Definition:
Given a product-related question and reviews as supporting documents, generate the answer to the question.
Data Processing:
This task only involves answerable and open-ended questions. To avoid data leakage, we use the Electronics and Home categories, which are different from the task, as the IND sources and the Cells category as the OOD source. Every data entry in the AmazonQA has a list of answers labeled with helpfulness levels. We choose the most helpful answer as the ground truth.
Evaluation Metrics:
The task is evaluated on precision, recall, and F1 of the BERTScoreΒ (Zhang etΒ al., 2019), and BLEURTΒ (Sellam etΒ al., 2020). BERTScore and BLEURT both evaluate the similarity of sentences by leveraging contextual embeddings from pre-trained models and are widely used in the NLP field. In this task, we use F1 BERTScore as the primary metric.
Example:
β’
Input:
β
Question: Can you add additional receiver with just one sensor? So one sensor picks up the the signal and sends it to two receiver.
β
[…]
β
β’
Output:
β
Yes...just be sure all have the same dip switch settings.
β
Note that the answer to the above question that is semantically similar to the output should be generated from the document.
Appendix B Data Statistics
The comprehensive and wide-ranging tasks in are critical for developing versatile e-commerce foundation models. includes products from 21 categories, such as Sports, Electronics, and Home. This broad coverage ensures that encompasses diverse product information, facilitating the creation of generalist e-commerce models.
Figure A1 shows the distribution of product categories for , , , , , and . The diverse categories in significantly enhance the effectiveness of . Note that the data of , , , and tasks lacks category information, thus their distributions are not shown.
Besides product categories, we also present the distributions of input lengths for each task, measured by word count, in Figure A2. For better clarity, we exclude very long inputs (those representing at most 1% of samples) in the , , , and tasks.
Due to limited computing resources, especially in academic settings, the current dataset has at most 10k training samples per task to fit within the model training capacity. However, scaling up is straightforward by following our data processing procedures detailed in Section 3 and Appendix A. For example, approximately 4.8 million high-quality samples could be further included in the task as outlined in Appendix A.1.1. To allow researchers and developers to customize to their specific needs, all data processing scripts are available at https://github.com/ninglab/eCeLLM/tree/main/data_processing.
Table A4 provides the number of samples at different stages: in the raw data, after data filtering and quality checks, and after sampling. The βAfter Filteringβ column shows the remaining data after applying quality checks. The βAfter Samplingβ column shows the data included in . Detailed statistics for the βAfter Samplingβ column, including training, validation, and testing samples for both in-domain and out-of-domain evaluation, are available in Table A3.
Even after filtering and quality checks, a substantial number of samples remain available for each task, making easily extendable and scalable. We believe can significantly benefit e-commerce foundation model research in various applications.
| Task | Raw Data | After Filtering | After Sampling |
| 15,177,901 | 4,767,579 | 13,000 | |
| 10,868,853 | 7,705,363 | 13,000 | |
| 2,600 | 2,528 | 2,528 | |
| 800,000 | 617,200 | 13,000 | |
| 31,013,687 | 1,946,393 | 13,000 | |
| 2,621,288 | 935,757 | 12,000 | |
| 2,621,288 | 935,757 | 12,000 | |
| 130,652 | 57,707 | 12,000 | |
| 220,541 | 184,814 | 13,000 | |
| 398,654 | 213,612 | 13,000 |
β’
The βAfter Samplingβ column indicates the data included in .
Appendix C Instructions
Each task has 6 instructions. One of them is the human-written seed instruction, 4 of them are GPT-4 generated diverse instructions, and the last one is the βunseenβ instruction.
C.1 Instructions on Attribute Value Extraction ()
Seed Instruction Given the title, description, feature, price, and brand of a product and a set of target attributes, extract the value of each target attribute from the product information. Output the extracted value and the corresponding source (e.g., title or feature) denoting where the value is extracted.
Generated Instruction 1 Extract the value of the target attribute from the given product information and output it along with the corresponding source.
Generated Instruction 2 Parse the product information to locate the target attribute, and then provide the extracted value of the target attribute and its source in the output, specifying None if the attribute is not present.
Generated Instruction 3 First, identify the target attributes from the provided list. Then, scan the product title, description, feature, and brand to extract the values associated with each target attribute. Finally, create a list of dictionaries, each containing the extracted attribute, its corresponding value, and the source where it was found.
Generated Instruction 4 Using the productβs title, description, features, price, and brand, identify and retrieve the values associated with a specified set of target attributes. Output the extracted values along with their respective sources (e.g., title or feature) indicating where each value was found.
Unseen Instruction Retrieve the value associated with the target attribute from the product information and specify the source (e.g., title, description, feature, or title) where the value was found.
C.2 Instructions on Product Relation Prediction ()
Seed Instruction Given the title of two products, predict if the two products are similar, if the two products will be purchased or viewed together. Answer only from the options.
Generated Instruction 1 Analyze the titles of Product 1 and Product 2 to determine if they are similar, if they will be purchased or viewed together, and choose the corresponding option.
Generated Instruction 2 Evaluate the titles of Product 1 and Product 2, then choose the option that best describes the relation between the two products.
Generated Instruction 3 Evaluate the titles of Product 1 and Product 2 to assess their similarity and whether they are likely to be purchased or viewed together. Then, select the appropriate option.
Generated Instruction 4 Predict whether two products are similar, whether two products are likely to be purchased or viewed together based on their titles. Choose your answer from the provided options.
Unseen Instruction Analyze the titles of Product 1 and Product 2 and select the option that indicates the relation of the two products.
C.3 Instructions on Product Matching ()
Seed Instruction Given the title, description, manufacturer, and price of two products, identify if they are the same product. Only output yes or no.
Generated Instruction 1 Analyze the title, description, manufacturer, and price between the two products below and generate an output of yes if the two products are the same, otherwise respond with no.
Generated Instruction 2 Check the details of the two products to see if they refer to the same product. Output only yes or no.
Generated Instruction 3 Based on the product information, predict if the two products are identical or not. Output yes if they are identical or no otherwise.
Generated Instruction 4 Compare the details of two given products to determine if they are identical. Output yes if they are identical or no otherwise.
Unseen Instruction Determine whether the two products are the same by comparing their title, description, manufacturer, and price, and provide a simple yes or no answer as the output.
C.4 Instructions on Sentiment Analysis ()
Seed Instruction Given the userβs review, identify the userβs sentiment from the listed options. Answer using one of the options.
Generated Instruction 1 Assess the userβs sentiment in the provided review and select the appropriate sentiment option from the list as the answer.
Generated Instruction 2 Determine the sentiment expressed by the user in her review from the provided choices, and respond by selecting one of the available options.
Generated Instruction 3 Carefully assess the userβs review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
Generated Instruction 4 Analyze the userβs review text and determine the overall sentiment expressed, then choose the corresponding sentiment option from the provided list (A: very positive, B: positive, C: neutral, D: negative, E: very negative) based on the identified sentiment.
Unseen Instruction Analyze the userβs review and determine the sentiment based on the listed options.
C.5 Instructions on Sequential Recommendation ()
Seed Instruction Given the products the user has purchased in history, rank the items in the listed options and output the item that the user is most likely to purchase next. Answer from one of the options.
Generated Instruction 1 Based on the userβs historical purchases, rank the items in options and predict the next product of the userβs interest from the provided options.
Generated Instruction 2 Rank the items in options and predict the userβs next purchase from the listed options by analyzing her historical purchases.
Generated Instruction 3 The userβs purchase history implies her preferences. Rank the items in the options based on the userβs preferences. Output the item that the user is most likely to purchase next from the options.
Generated Instruction 4 Rank items in listed options based on the userβs purchase history to determine the item that the user is most likely to purchase next. Output the item with the highest likelihood of being the next purchase.
Unseen Instruction Estimate the userβs intent based on the userβs purchase history, and predict the next product that the user is most likely to purchase from the given options.
C.6 Instructions on Multi-class Product Classification ()
Seed Instruction What is the relevance between the query and the product title below? Answer from one of the options.
Generated Instruction 1 Analyze the query and product title to determine the relevance between the query and product, and select the appropriate option from the provided options.
Generated Instruction 2 Evaluate the relevance between the query and product title, and choose the most accurate option from the given options.
Generated Instruction 3 Analyze the query and product title to assess the level of relevance between them, and then output the corresponding option that best describes this relevance.
Generated Instruction 4 Determine the relevance between the query and the product title provided, and select your response from one of the available options.
Unseen Instruction Compare the query and the product title to determine if the product fully meets the query specifications. Choose the option that best describes the relevance between them.
C.7 Instructions on Product Substitute Identification ()
Seed Instruction Given a query and a product, identify if the product is somewhat relevant to the query. It fails to fulfill some aspects of the query but the product can be used as a functional substitute. Only output yes or no.
Generated Instruction 1 Answer yes if the product is a substitute for the query and no otherwise.
Generated Instruction 2 Please respond with yes if the product is a suitable substitute for the query, and no if it is not.
Generated Instruction 3 Check if a product can function as a substitute for a given query, even if it doesnβt fully meet all requirements. Output yes if it can or no otherwise.
Generated Instruction 4 Assess the relevance of a product to a given query by determining if it can function as a substitute, despite not fully meeting certain aspects of the query. Provide a binary output of yes or no based on this evaluation.
Unseen Instruction Assess whether the product is a substitute for the query and provide a yes or no response.
C.8 Instructions on Query Product Ranking ()
Seed Instruction Given a query and a list of products denoted as A, B, C, β¦ with their titles, rank the products according to their relevance to the query. Output only a ranked list in which the most relevant product is at the top of the list.
Generated Instruction 1 Evaluate each product title in the given list, assess its relevance to the given query, and then arrange the products in descending order of relevance, with the most relevant product at the top of the ranked list.
Generated Instruction 2 Rank the products A, B, C, β¦ based on their relevance to the provided query, and produce a ranked list with the most relevant product positioned at the top of the list.
Generated Instruction 3 Analyze the query and each product title. Sort the products in descending order based on their relevance to the query. The most relevant product should be at the top of the list, and output the ranked list.
Generated Instruction 4 Evaluate the relevance of each product title in the input to the given query, and then sort the products in descending order of relevance, placing the most relevant product at the top of the ranked list.
Unseen Instruction Evaluate the query against each productβs title, determine the relevance between the query and the product, and organize the products in descending order of relevance, ensuring that the product with the highest relevance is positioned at the top of the list.
C.9 Instructions on Answerability Prediction ()
Seed Instruction Given a question and the related document, predict if the question is answerable based on the information provided in the document. Output only yes or no.
Generated Instruction 1 Evaluate the answerability of a question by analyzing the related document, outputting yes if the document contains information addressing the question, and no otherwise.
Generated Instruction 2 Analyze a question and its supporting document. Predicting answerability based on the information provided in the document. Output yes if the document contains relevant information to answer the question, otherwise output no.
Generated Instruction 3 Given a question and its related document, determine if the question is answerable by analyzing the information in the document. Output yes if the document addresses the question, or no otherwise.
Generated Instruction 4 Output yes if the supporting document can answer the given question. Otherwise, output no.
Unseen Instruction Predict whether it is possible to answer the given question using the supporting document, and output a yes or no response.
C.10 Instructions on Answer Generation ()
Seed Instruction Given a question and the related document, and generate the answer to the question based on the information provided in the document.
Generated Instruction 1 Generate an answer to the question by utilizing the information contained in the document.
Generated Instruction 2 Extract information from the supporting document to answer the given question.
Generated Instruction 3 Answer the given question using the supporting document.
Generated Instruction 4 Answer the given question by extracting information from the supporting document.
Unseen Instruction Utilize the information provided in the supporting document to generate an answer to the given question.
Appendix D Training and Evaluation Details
The training details of general-purpose and e-commerce baseline LLMs are shown in TableΒ A5. The general-purpose LLMs undergo the 1-shot evaluation, which measures the sample with one in-context example and one test case. For each task, the 1-shot evaluation dataset is composed of all 1K test samples and the same amount of training samples (except for , which uses all 253 test samples and 253 training samples) from its own test and training sets. The e-commerce LLM undergoes both 1-shot and 0-shot evaluation.
| General-purpose LLMs | URL | Accessibility |
| GPT-4 Turbo | https://platform.openai.com/ | API |
| Gemini Pro | https://ai.google.dev/tutorials/python_quickstart | API |
| Claude 2.1 | https://docs.anthropic.com/claude/reference/getting-started-with-the-api | API |
| Llama-2 13B-chat | https://huggingface.co/meta-llama/Llama-2-13b-chat-hf | Checkpoint |
| Mistral-7B-Instruct-v0.2 | https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 | Checkpoint |
| EcomGPT | https://github.com/Alibaba-NLP/EcomGPT | Checkpoint |
| Tasks | Baseline Models | Model Source | Mode | Parameters |
| SUOpenTagΒ (Xu etΒ al., 2019) | https://github.com/hackerxiaobai/OpenTag_2019 | training | epoch: 100 | |
| AVEQAΒ (Wang etΒ al., 2020) | https://github.com/Zinc-30/aveqa | training | epoch: 89; batch size: 16 | |
| RGCNΒ (Schlichtkrull etΒ al., 2018) | https://github.com/JinheonBaek/RGCN | training | epochs: 500K | |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 | |
| BERTweetΒ (Nguyen etΒ al., 2020) | https://huggingface.co/finiteautomata/bertweet-base-sentiment-analysis | training | epoch: 3; Β Β Β Β batch size: 8 | |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 | |
| P5Β (Geng etΒ al., 2022) | https://github.com/jeykigung/P5 | evaluate on checkpoint | checkpoint: toy base | |
| gSASRecΒ (Petrov & Macdonald, 2023) | https://github.com/asash/gSASRec-pytorch | training | sequence length: 50 | |
| RecformerΒ (Li etΒ al., 2023) | https://github.com/AaronHeee/RecFormer | fine-tuning | epoch: 16; Β Β Β Β batch size 8 | |
| GPT-4 TurboΒ (OpenAI, 2023) | https://platform.openai.com/ | 1-shot Β Β Β learning | prompt template inΒ E.2 | |
| , , , , | BERTΒ (Devlin etΒ al., 2019) | https://huggingface.co/bert-base-multilingual-cased | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 |
| DeBERTaV3Β (He etΒ al., 2022) | https://huggingface.co/microsoft/deberta-v3-base | fine-tuning | epoch: 3; Β Β Β Β batch size: 8 |
β’
In this table, βtrainingβ means training from scratch, βevaluate on checkpointβ means that we evaluate using the checkpoint provided by the link, βfine-tuningβ means that we fine-tune the checkpoint provided by the link on specific tasks, and β1-shot learningβ indicates that we directly use the model checkpoint from the link, and prompt the model with one in-context example.
The training details of SoTA task-specific models are presented in Table A6. For each task, we train and test on the SoTA task-specific models of each task using its own training, validation, and IND test sets (i.e., task-specific). For the tasks with the OOD test set, we save the trained model and test them on the OOD test set. For SUOpenTagΒ (Xu etΒ al., 2019) and AVEQAΒ (Wang etΒ al., 2020), we evaluate the task according to the equationΒ 1. Note that we do not consider MAVEQAΒ (Yang etΒ al., 2022) as the baseline of , as it has demonstrated similar performance to AVEQA. For RGCNΒ (Schlichtkrull etΒ al., 2018), we construct a product-product graph, in which the nodes are products, and the edges are specified by the relations between products (e.g., similar products). We initialize the node embeddings in the RGCN using the embeddings of product titles generated from BERTΒ (Devlin etΒ al., 2019). The BERT, DeBERTa, and BERTweet baselines are implemented through sentence-transformerΒ (Reimers & Gurevych, 2019). For P5Β (Geng etΒ al., 2022), we conduct zero-shot evaluations on beauty, sports, and toys base checkpoints, and report the best result (toy base). Similar to the RGCN, we use the BERT-generated embeddings of product titles to initialize item embeddings in the gSASRec. For both gSASRec and RecformerΒ (Li etΒ al., 2023), we evaluate the results of the next product of interest within the candidate list as detailed in AppendixΒ A.2.2
Appendix E Prompt Templates
The following prompts are used for evaluating general-purpose and e-commerce LLMs. For Claude 2.1, the prompt begins with {HUMAN PROMPT} and ends with {AI PROMPT}. For Mistral-7B Instruct-v0.2, the prompt is wrapped with β[INST]β.
E.1 Prompt Templates for Zero-shot Evaluation
Prompt with Options
β’
System prompt: Below is an instruction that describes a task. Write a response that appropriately completes the request.
β’
Instruction: {instruction}
β
input: {input}
β
options: {options}
β
response:
β
Prompt without Options
β’
System prompt: Below is an instruction that describes a task. Write a response that appropriately completes the request.
β’
Instruction: {instruction}
β
input: {input}
β
response:
β
E.2 Prompt Templates for One-shot Evaluation
Prompt with Options
β’
System prompt: Below is an instruction that describes a task, paired with an example that provides further context for the task.
β’
Instruction: {instruction}
β’
Example:
β
input: {example.input}
β
options: {example.options}
β
response: {example.response}
β
β’
Now write a response that appropriately completes the following example.
β
input: {input}
β
options: {options}
β
response:
β
Prompt without Options
β’
System prompt: Below is an instruction that describes a task, paired with an example that provides further context for the task.
β’
Instruction: {instruction}
β’
Example:
β
input: {example.input}
β
response: {example.response}
β
β’
Now write a response that appropriately completes the following example.
β
input: {input}
β
response:
β
Appendix F Analysis on Base Models
| Model | Base Model | ||||||||||
| F1* | Macro F1 | F1 | Macro F1 | HR@1 | Accuracy | F1 | NDCG | F1 | F | ||
| Flan-T5 XXL | 0.447 | 0.522 | 0.995 | 0.628 | 0.512 | 0.680 | 0.376 | 0.885 | 0.836 | 0.844 | |
| Llama-2 13B-chat | 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 | |
| Llama-2 7B-chat | 0.577 | 0.562 | 0.995 | 0.613 | 0.503 | 0.695 | 0.444 | 0.864 | 0.859 | 0.839 | |
| Mistral-7B Instruct-v0.2 | 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 | |
| Flan-T5 XL | 0.376 | 0.511 | 0.995 | 0.648 | 0.463 | 0.663 | 0.042 | 0.868 | 0.827 | 0.843 | |
| Phi-2 | 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 |
β’
In this table, βBase Modelβ presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
We compare the 6 base models considered for the series of models (i.e., , , and ) and show the performance comparison in TableΒ A7. From TableΒ A7, we observe that Llama-2 13B-chat is the best-performing base model for . The instruction-tuned Llama-2 13B-chat model demonstrates considerable improvement compared to instruction-tuned Flan-T5 XXL on 7 out of the 10 tasks. We also observe that Mistral-7B Instruct-v0.2 and Phi-2 are the best-performing base models for and , respectively. Particularly, with instruction tuning, Mistral-7B Instruct-v0.2 achieves a notable average improvement of 2.2% over Llama-2 7B-chat across the 10 tasks. Similarly, Instruction-tuned Phi-2 also outperforms Flan-T5 XL on 6 out of the 10 tasks and achieves similar performance with Flan-T5 XL on the rest 4 tasks. The variations of trainable size and focused aspect contribute to the distinct inherent capabilities of the base models, which play a crucial role in adapting LLMs to e-commerce scenarios.
Appendix G Complete Experimental Results
The complete results for both IND and OOD evaluations for tasks (1) attribute value extraction (), (2) product relation prediction (), (3) product matching (), (4) sentiment analysis (), (5) sequential recommendation (), (6) multi-class product classification (), (7) product substitute identification (), (8) query product ranking (), (9) answerability prediction (), and (10) answer generation () are presented in TableΒ A8, Β A9, Β A10, Β A11, Β A16, Β A12, Β A13, Β A16, Β A14, andΒ A15, respectively. Overall, these tables show that models fine-tuned on outperform the general-purpose LLMs and SoTA task-specific models in the IND test. Meanwhile, the exhibits good generalizability to the OOD data. Because of the high-quality , achieves remarkable performance with different base models. Note that #failed in tables represents the number of failure cases for which we cannot extract meaningful results from the model output. These failure cases are counted as wrong predictions when calculating the metrics. NPR refers to the negative prediction rate.
Regarding the hallucination issue, we observe very limited hallucination from . As shown in the #failed column of TableΒ A8 and Β A16, in and tasks, could slightly suffer from hallucination and does not output results in the desired format for several testing samples. For the other tasks except for the task, as shown in TableΒ A9-A14, is robust to hallucination and could always output options in a desired format. For the task, we randomly sample 50 answers generated by and do not observe hallucination with manual checks.
| Model | IND | OOD | |||||||||
| Recall* | Precision* | F1* | #failed | Recall* | Precision* | F1* | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.422 | 0.598 | 0.495 | 6 | 0.317 | 0.529 | 0.397 | 1 | ||
| Gemini Pro | 0.318 | 0.523 | 0.396 | 4 | 0.203 | 0.426 | 0.275 | 6 | |||
| Claude 2.1 | 0.310 | 0.494 | 0.381 | 59 | 0.312 | 0.600 | 0.410 | 66 | |||
| Llama-2 13B-chat | 0.002 | 0.002 | 0.002 | 0 | 0.000 | 0.000 | 0.000 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.321 | 0.435 | 0.369 | 69 | 0.217 | 0.337 | 0.264 | 52 | |||
| E-commerce LLM | EcomGPT | 0.000 | 0.000 | 0.000 | 905 | 0.001 | 0.042 | 0.001 | 869 | ||
| SoTA task-specific model | SUOpenTag | 0.603 | 0.500 | 0.546 | 0 | 0.124 | 0.173 | 0.144 | 0 | ||
| AVEQA | 0.425 | 0.491 | 0.456 | 0 | 0.283 | 0.257 | 0.269 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.298 | 0.519 | 0.378 | 7 | 0.362 | 0.701 | 0.477 | 0 | ||
| Llama-2 13B-chat | 0.544 | 0.666 | 0.599 | 3 | 0.448 | 0.613 | 0.518 | 2 | |||
| Llama-2 7B-chat | 0.531 | 0.660 | 0.588 | 1 | 0.323 | 0.499 | 0.392 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.720 | 0.799 | 0.757 | 5 | 0.374 | 0.544 | 0.443 | 0 | |||
| Flan-T5 XL | 0.258 | 0.449 | 0.328 | 7 | 0.276 | 0.538 | 0.365 | 3 | |||
| Phi-2 | 0.304 | 0.570 | 0.397 | 265 | 0.288 | 0.488 | 0.362 | 0 | |||
| Generalist | Flan-T5 XXL | 0.353 | 0.611 | 0.447 | 2 | 0.360 | 0.699 | 0.476 | 0 | ||
| Llama-2 13B-chat | 0.530 | 0.646 | 0.582 | 1 | 0.276 | 0.425 | 0.335 | 0 | |||
| Llama-2 7B-chat | 0.514 | 0.641 | 0.571 | 4 | 0.236 | 0.392 | 0.294 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.612 | 0.722 | 0.662 | 0 | 0.304 | 0.463 | 0.367 | 0 | |||
| Flan-T5 XL | 0.297 | 0.514 | 0.376 | 1 | 0.267 | 0.518 | 0.352 | 1 | |||
| Phi-2 | 0.455 | 0.578 | 0.509 | 0 | 0.237 | 0.417 | 0.302 | 0 | |||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. Recall*, Precision* and F1* are defined as equationΒ 1 in AppendixΒ A.1.1, and #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| Accuracy | M-Rec | M-Pre | M-F1 | #failed | Accuracy | M-Rec | M-Pre | M-F1 | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.384 | 0.487 | 0.381 | 0.326 | 0 | 0.488 | 0.496 | 0.392 | 0.392 | 0 | ||
| Gemini Pro | 0.128 | 0.385 | 0.352 | 0.136 | 1 | 0.147 | 0.359 | 0.390 | 0.123 | 0 | |||
| Claude 2.1 | 0.508 | 0.347 | 0.344 | 0.275 | 10 | 0.362 | 0.394 | 0.400 | 0.277 | 4 | |||
| Llama-2 13B-chat | 0.473 | 0.333 | 0.333 | 0.333 | 0 | 0.419 | 0.338 | 0.339 | 0.324 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.442 | 0.323 | 0.325 | 0.324 | 0 | 0.422 | 0.338 | 0.351 | 0.327 | 0 | |||
| E-commerce LLM | EcomGPT | 0.147 | 0.101 | 0.101 | 0.091 | 444 | 0.125 | 0.125 | 0.092 | 0.096 | 455 | ||
| SoTA task-specific model | DeBERTaV3 | 0.762 | 0.575 | 0.620 | 0.588 | 0 | 0.658 | 0.514 | 0.570 | 0.507 | 0 | ||
| RGCN | 0.615 | 0.665 | 0.637 | 0.506 | 0 | 0.576 | 0.373 | 0.372 | 0.356 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.754 | 0.516 | 0.511 | 0.508 | 0 | 0.663 | 0.506 | 0.468 | 0.466 | 0 | ||
| Llama-2 13B-chat | 0.769 | 0.530 | 0.517 | 0.521 | 0 | 0.690 | 0.520 | 0.472 | 0.483 | 0 | |||
| Llama-2 7B-chat | 0.774 | 0.541 | 0.628 | 0.537 | 0 | 0.695 | 0.526 | 0.803 | 0.498 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.782 | 0.547 | 0.689 | 0.543 | 0 | 0.711 | 0.532 | 0.808 | 0.502 | 0 | |||
| Flan-T5 XL | 0.704 | 0.467 | 0.496 | 0.460 | 0 | 0.592 | 0.471 | 0.625 | 0.427 | 0 | |||
| Phi-2 | 0.584 | 0.372 | 0.379 | 0.348 | 0 | 0.406 | 0.349 | 0.334 | 0.251 | 0 | |||
| Generalist | Flan-T5 XXL | 0.769 | 0.531 | 0.517 | 0.522 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | ||
| Llama-2 13B-chat | 0.775 | 0.599 | 0.635 | 0.611 | 0 | 0.726 | 0.564 | 0.611 | 0.558 | 0 | |||
| Llama-2 7B-chat | 0.797 | 0.586 | 0.661 | 0.595 | 0 | 0.703 | 0.533 | 0.648 | 0.499 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.788 | 0.555 | 0.644 | 0.558 | 0 | 0.707 | 0.537 | 0.596 | 0.502 | 0 | |||
| Flan-T5 XL | 0.757 | 0.517 | 0.515 | 0.511 | 0 | 0.678 | 0.521 | 0.587 | 0.489 | 0 | |||
| Phi-2 | 0.747 | 0.524 | 0.552 | 0.518 | 0 | 0.710 | 0.541 | 0.611 | 0.520 | 0 | |||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βM-Recβ, βM-Preβ, and βM-F1β represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
| Model | IND | ||||||||
| Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.826 | 0.604 | 1.000 | 0.753 | 1.000 | 0.763 | 0 | |
| Gemini Pro | 0.897 | 0.766 | 1.000 | 0.867 | 1.000 | 0.845 | 0 | ||
| Claude 2.1 | 0.711 | 0.360 | 0.952 | 0.523 | 0.986 | 0.664 | 1 | ||
| Llama-2 13B-chat | 0.474 | 0.459 | 0.411 | 0.434 | 0.486 | 0.535 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.755 | 0.441 | 1.000 | 0.613 | 1.000 | 0.696 | 0 | ||
| E-commerce LLM | EcomGPT | 0.648 | 0.739 | 0.577 | 0.648 | 0.577 | 0.739 | 0 | |
| SoTA task-specific model | BERT | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | |
| DeBERTaV3 | 0.577 | 1.000 | 0.509 | 0.675 | 0.246 | 1.000 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | |
| Llama-2 13B-chat | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | ||
| Llama-2 7B-chat | 0.992 | 0.991 | 0.991 | 0.991 | 0.993 | 0.993 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.988 | 0.991 | 0.982 | 0.987 | 0.986 | 0.993 | 0 | ||
| Flan-T5 XL | 0.960 | 0.910 | 1.000 | 0.953 | 1.000 | 0.934 | 0 | ||
| Phi-2 | 0.992 | 0.991 | 0.991 | 0.991 | 0.993 | 0.993 | 0 | ||
| Generalist | Flan-T5 XXL | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | |
| Llama-2 13B-chat | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | ||
| Llama-2 7B-chat | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | ||
| Flan-T5 XL | 0.996 | 0.991 | 1.000 | 0.995 | 1.000 | 0.993 | 0 | ||
| Phi-2 | 0.992 | 0.991 | 0.991 | 0.991 | 0.993 | 0.993 | 0 | ||
β’
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| Acc | M-Rec | M-Pre | M-F1 | #failed | Acc | M-Rec | M-Pre | M-F1 | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.595 | 0.575 | 0.544 | 0.516 | 0 | 0.556 | 0.586 | 0.544 | 0.510 | 0 | ||
| Gemini Pro | 0.609 | 0.521 | 0.453 | 0.470 | 2 | 0.572 | 0.511 | 0.444 | 0.454 | 1 | |||
| Claude 2.1 | 0.375 | 0.510 | 0.474 | 0.415 | 2 | 0.328 | 0.466 | 0.447 | 0.369 | 1 | |||
| Llama-2 13B-chat | 0.406 | 0.188 | 0.191 | 0.188 | 0 | 0.384 | 0.179 | 0.180 | 0.178 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.633 | 0.532 | 0.551 | 0.470 | 0 | 0.594 | 0.531 | 0.494 | 0.438 | 0 | |||
| E-commerce LLM | EcomGPT | 0.191 | 0.362 | 0.341 | 0.188 | 6 | 0.196 | 0.375 | 0.336 | 0.178 | 13 | ||
| SoTA task-specific model | BERTweet | 0.733 | 0.503 | 0.530 | 0.511 | 0 | 0.729 | 0.507 | 0.524 | 0.513 | 0 | ||
| DeBERTaV3 | 0.768 | 0.567 | 0.607 | 0.573 | 0 | 0.764 | 0.565 | 0.591 | 0.567 | 0 | |||
| P5 | 0.611 | 0.199 | 0.157 | 0.156 | 0 | 0.620 | 0.200 | 0.124 | 0.153 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.783 | 0.619 | 0.618 | 0.612 | 0 | 0.770 | 0.604 | 0.601 | 0.600 | 0 | ||
| Llama-2 13B-chat | 0.791 | 0.616 | 0.641 | 0.616 | 0 | 0.781 | 0.627 | 0.645 | 0.629 | 0 | |||
| Llama-2 7B-chat | 0.790 | 0.620 | 0.652 | 0.634 | 0 | 0.769 | 0.583 | 0.599 | 0.589 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.801 | 0.643 | 0.676 | 0.655 | 0 | 0.789 | 0.619 | 0.650 | 0.632 | 0 | |||
| Flan-T5 XL | 0.771 | 0.645 | 0.638 | 0.620 | 0 | 0.743 | 0.594 | 0.592 | 0.582 | 0 | |||
| Phi-2 | 0.779 | 0.611 | 0.618 | 0.608 | 0 | 0.754 | 0.576 | 0.594 | 0.583 | 0 | |||
| Generalist | Flan-T5 XXL | 0.797 | 0.629 | 0.646 | 0.628 | 0 | 0.787 | 0.619 | 0.624 | 0.619 | 0 | ||
| Llama-2 13B-chat | 0.796 | 0.641 | 0.661 | 0.648 | 0 | 0.785 | 0.621 | 0.638 | 0.629 | 0 | |||
| Llama-2 7B-chat | 0.768 | 0.579 | 0.589 | 0.580 | 0 | 0.776 | 0.599 | 0.626 | 0.606 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.781 | 0.630 | 0.654 | 0.639 | 0 | 0.784 | 0.630 | 0.653 | 0.640 | 0 | |||
| Flan-T5 XL | 0.782 | 0.654 | 0.655 | 0.648 | 0 | 0.753 | 0.604 | 0.598 | 0.598 | 0 | |||
| Phi-2 | 0.780 | 0.588 | 0.619 | 0.596 | 0 | 0.758 | 0.552 | 0.590 | 0.565 | 0 | |||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βAccβ, βM-Recβ, βM-Preβ, and βM-F1β represents accuracy, macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
| Model | IND | ||||||
| Accuracy | M-Rec | M-Pre | M-F1 | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.611 | 0.527 | 0.540 | 0.487 | 0 | |
| Gemini Pro | 0.584 | 0.471 | 0.414 | 0.425 | 2 | ||
| Claude 2.1 | 0.655 | 0.464 | 0.419 | 0.435 | 13 | ||
| Llama-2 13B-chat | 0.504 | 0.250 | 0.251 | 0.250 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.529 | 0.395 | 0.384 | 0.365 | 0 | ||
| E-commerce LLM | EcomGPT | 0.540 | 0.265 | 0.218 | 0.223 | 2 | |
| SoTA task-specific model | BERT | 0.661 | 0.381 | 0.423 | 0.393 | 0 | |
| DeBERTaV3 | 0.703 | 0.436 | 0.472 | 0.448 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.666 | 0.438 | 0.412 | 0.346 | 0 | |
| Llama-2 13B-chat | 0.655 | 0.399 | 0.410 | 0.349 | 0 | ||
| Llama-2 7B-chat | 0.659 | 0.399 | 0.531 | 0.330 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.681 | 0.406 | 0.423 | 0.387 | 0 | ||
| Flan-T5 XL | 0.648 | 0.425 | 0.361 | 0.327 | 0 | ||
| Phi-2 | 0.646 | 0.387 | 0.316 | 0.321 | 0 | ||
| Generalist | Flan-T5 XXL | 0.680 | 0.431 | 0.416 | 0.364 | 0 | |
| Llama-2 13B-chat | 0.684 | 0.440 | 0.435 | 0.414 | 0 | ||
| Llama-2 7B-chat | 0.679 | 0.427 | 0.434 | 0.398 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.696 | 0.450 | 0.456 | 0.443 | 0 | ||
| Flan-T5 XL | 0.663 | 0.395 | 0.533 | 0.332 | 0 | ||
| Phi-2 | 0.650 | 0.397 | 0.410 | 0.335 | 0 | ||
β’
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βM-Recβ, βM-Preβ, and βM-F1β represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
| Model | IND | ||||||||
| Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.289 | 0.374 | 0.132 | 0.195 | 0.264 | 0.585 | 0 | |
| Gemini Pro | 0.296 | 0.504 | 0.164 | 0.248 | 0.234 | 0.612 | 0 | ||
| Claude 2.1 | 0.291 | 0.578 | 0.179 | 0.273 | 0.205 | 0.620 | 1 | ||
| Llama-2 13B-chat | 0.649 | 0.257 | 0.247 | 0.252 | 0.766 | 0.775 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.361 | 0.609 | 0.203 | 0.305 | 0.287 | 0.711 | 0 | ||
| E-commerce LLM | EcomGPT | 0.630 | 0.165 | 0.176 | 0.170 | 0.769 | 0.755 | 13 | |
| SoTA task-specific model | BERT | 0.761 | 0.330 | 0.472 | 0.389 | 0.890 | 0.816 | 0 | |
| DeBERTaV3 | 0.769 | 0.000 | 0.000 | 0.000 | 0.999 | 0.770 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.766 | 0.013 | 0.300 | 0.025 | 0.991 | 0.771 | 0 | |
| Llama-2 13B-chat | 0.770 | 0.000 | 0.000 | 0.000 | 1.000 | 0.770 | 0 | ||
| Llama-2 7B-chat | 0.770 | 0.017 | 0.500 | 0.034 | 0.995 | 0.772 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.770 | 0.000 | 0.000 | 0.000 | 1.000 | 0.770 | 0 | ||
| Flan-T5 XL | 0.768 | 0.000 | 0.000 | 0.000 | 0.997 | 0.770 | 0 | ||
| Phi-2 | 0.770 | 0.000 | 0.000 | 0.000 | 1.000 | 0.770 | 0 | ||
| Generalist | Flan-T5 XXL | 0.771 | 0.300 | 0.504 | 0.376 | 0.912 | 0.813 | 0 | |
| Llama-2 13B-chat | 0.795 | 0.448 | 0.569 | 0.501 | 0.899 | 0.845 | 0 | ||
| Llama-2 7B-chat | 0.781 | 0.283 | 0.546 | 0.372 | 0.930 | 0.813 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.790 | 0.200 | 0.639 | 0.305 | 0.966 | 0.802 | 0 | ||
| Flan-T5 XL | 0.773 | 0.022 | 0.714 | 0.042 | 0.997 | 0.773 | 0 | ||
| Phi-2 | 0.777 | 0.313 | 0.526 | 0.392 | 0.916 | 0.817 | 0 | ||
β’
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||||||
| Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.623 | 0.550 | 0.791 | 0.649 | 0.749 | 0.491 | 5 | 0.641 | 0.555 | 0.876 | 0.680 | 0.828 | 0.460 | 2 | ||
| Gemini Pro | 0.542 | 0.371 | 0.797 | 0.506 | 0.837 | 0.435 | 13 | 0.543 | 0.410 | 0.844 | 0.552 | 0.834 | 0.393 | 6 | |||
| Claude 2.1 | 0.424 | 0.177 | 0.671 | 0.280 | 0.850 | 0.375 | 90 | 0.384 | 0.146 | 0.769 | 0.245 | 0.904 | 0.326 | 73 | |||
| Llama-2 13B-chat | 0.534 | 0.608 | 0.638 | 0.623 | 0.406 | 0.375 | 0 | 0.541 | 0.605 | 0.688 | 0.644 | 0.401 | 0.317 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.522 | 0.539 | 0.647 | 0.588 | 0.493 | 0.383 | 120 | 0.537 | 0.523 | 0.725 | 0.608 | 0.567 | 0.352 | 126 | |||
| E-commerce LLM | EcomGPT | 0.318 | 0.051 | 0.283 | 0.086 | 0.779 | 0.322 | 254 | 0.286 | 0.085 | 0.403 | 0.140 | 0.726 | 0.266 | 245 | ||
| SoTA task-specific model | BERT | 0.749 | 0.970 | 0.726 | 0.830 | 0.368 | 0.877 | 0 | 0.803 | 0.831 | 0.876 | 0.853 | 0.742 | 0.668 | 0 | ||
| DeBERTaV3 | 0.504 | 0.310 | 0.769 | 0.441 | 0.839 | 0.413 | 0 | 0.501 | 0.345 | 0.826 | 0.487 | 0.841 | 0.370 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.749 | 0.875 | 0.763 | 0.815 | 0.531 | 0.712 | 0 | 0.799 | 0.889 | 0.830 | 0.859 | 0.602 | 0.713 | 0 | ||
| Llama-2 13B-chat | 0.801 | 0.919 | 0.797 | 0.854 | 0.597 | 0.811 | 0 | 0.838 | 0.924 | 0.852 | 0.887 | 0.650 | 0.797 | 0 | |||
| Llama-2 7B-chat | 0.741 | 0.889 | 0.749 | 0.813 | 0.485 | 0.718 | 0 | 0.761 | 0.864 | 0.802 | 0.832 | 0.535 | 0.644 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.821 | 0.896 | 0.834 | 0.864 | 0.692 | 0.794 | 0 | 0.835 | 0.891 | 0.872 | 0.881 | 0.713 | 0.749 | 0 | |||
| Flan-T5 XL | 0.693 | 0.684 | 0.802 | 0.738 | 0.708 | 0.565 | 0 | 0.707 | 0.682 | 0.862 | 0.762 | 0.761 | 0.523 | 0 | |||
| Phi-2 | 0.765 | 0.942 | 0.751 | 0.835 | 0.460 | 0.820 | 0 | 0.781 | 0.950 | 0.779 | 0.856 | 0.411 | 0.791 | 0 | |||
| Generalist | Flan-T5 XXL | 0.774 | 0.910 | 0.773 | 0.836 | 0.540 | 0.776 | 0 | 0.814 | 0.914 | 0.832 | 0.871 | 0.596 | 0.760 | 0 | ||
| Llama-2 13B-chat | 0.808 | 0.864 | 0.838 | 0.851 | 0.711 | 0.752 | 0 | 0.818 | 0.862 | 0.872 | 0.867 | 0.723 | 0.705 | 0 | |||
| Llama-2 7B-chat | 0.817 | 0.921 | 0.814 | 0.864 | 0.638 | 0.824 | 0 | 0.836 | 0.931 | 0.845 | 0.886 | 0.627 | 0.807 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.797 | 0.880 | 0.814 | 0.846 | 0.654 | 0.759 | 0 | 0.832 | 0.885 | 0.872 | 0.878 | 0.717 | 0.740 | 0 | |||
| Flan-T5 XL | 0.765 | 0.888 | 0.774 | 0.827 | 0.553 | 0.741 | 0 | 0.819 | 0.891 | 0.852 | 0.871 | 0.662 | 0.735 | 0 | |||
| Phi-2 | 0.794 | 0.897 | 0.801 | 0.846 | 0.616 | 0.777 | 0 | 0.823 | 0.937 | 0.828 | 0.879 | 0.573 | 0.807 | 0 | |||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| R | P | F | BLEURT | #failed | R | P | F | BLEURT | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | ||
| Gemini Pro | 0.844 | 0.867 | 0.855 | 0.269 | 0 | 0.847 | 0.866 | 0.856 | 0.264 | 0 | |||
| Claude 2.1 | 0.848 | 0.835 | 0.841 | 0.314 | 0 | 0.851 | 0.833 | 0.842 | 0.325 | 0 | |||
| Llama-2 13B-chat | 0.845 | 0.780 | 0.811 | 0.261 | 0 | 0.845 | 0.775 | 0.808 | 0.260 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.850 | 0.856 | 0.853 | 0.288 | 0 | 0.852 | 0.851 | 0.851 | 0.290 | 0 | |||
| E-commerce LLM | EcomGPT | 0.675 | 0.665 | 0.669 | 0.290 | 0 | 0.729 | 0.716 | 0.722 | 0.296 | 0 | ||
| SoTA task-specific model | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.822 | 0.864 | 0.842 | 0.310 | 0 | 0.824 | 0.865 | 0.843 | 0.302 | 0 | ||
| Llama-2 13B-chat | 0.824 | 0.861 | 0.841 | 0.309 | 0 | 0.821 | 0.860 | 0.840 | 0.289 | 0 | |||
| Llama-2 7B-chat | 0.822 | 0.861 | 0.841 | 0.301 | 0 | 0.820 | 0.861 | 0.840 | 0.289 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.823 | 0.860 | 0.841 | 0.310 | 0 | 0.823 | 0.861 | 0.842 | 0.298 | 0 | |||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.320 | 0 | 0.824 | 0.864 | 0.843 | 0.307 | 0 | |||
| Phi-2 | 0.817 | 0.855 | 0.835 | 0.283 | 0 | 0.817 | 0.856 | 0.835 | 0.270 | 0 | |||
| Generalist | Flan-T5 XXL | 0.824 | 0.865 | 0.844 | 0.224 | 0 | 0.823 | 0.864 | 0.843 | 0.206 | 0 | ||
| Llama-2 13B-chat | 0.823 | 0.861 | 0.841 | 0.215 | 0 | 0.822 | 0.861 | 0.841 | 0.195 | 0 | |||
| Llama-2 7B-chat | 0.822 | 0.860 | 0.840 | 0.208 | 0 | 0.819 | 0.859 | 0.838 | 0.188 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.822 | 0.864 | 0.842 | 0.213 | 0 | 0.821 | 0.862 | 0.840 | 0.194 | 0 | |||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.227 | 0 | 0.824 | 0.865 | 0.844 | 0.211 | 0 | |||
| Phi-2 | 0.823 | 0.861 | 0.842 | 0.222 | 0 | 0.821 | 0.859 | 0.840 | 0.198 | 0 | |||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics βRβ, βPβ, βFβ and βBLEURTβ are detailed in TableΒ A1. On each task, the best performance is in bold. Note that we use GPT-4 Turbo as the SoTA task-specific model in this task.
| Model | ||||||||||
| IND | OOD | IND | ||||||||
| HR@1 | #failed | HR@1 | #failed | NDCG | #failed | |||||
| General-purpose LLMs | GPT-4 Turbo | 0.387 | 0 | 0.198 | 0 | 0.875 | 14 | |||
| Gemini Pro | 0.269 | 2 | 0.116 | 3 | 0.821 | 52 | ||||
| Claude 2.1 | 0.066 | 34 | 0.036 | 42 | 0.821 | 26 | ||||
| Llama-2 13B-chat | 0.056 | 0 | 0.050 | 0 | 0.815 | 0 | ||||
| Mistral-7B-Instruct-v0.2 | 0.164 | 1 | 0.108 | 0 | 0.842 | 4 | ||||
| E-commerce LLM | EcomGPT | 0.042 | 344 | 0.023 | 391 | 0.000 | 1000 | |||
| SoTA task-specific model | gSASRec / BERT | 0.249 | 0 | 0.065 | 0 | 0.839 | 0 | |||
| Recformer / DeBERTaV3 | 0.265 | 0 | 0.081 | 0 | 0.859 | 0 | ||||
| Task-specific | Flan-T5 XXL | 0.467 | 0 | 0.252 | 0 | 0.881 | 0 | |||
| Llama-2 13B-chat | 0.518 | 0 | 0.263 | 0 | 0.879 | 0 | ||||
| Llama-2 7B-chat | 0.517 | 0 | 0.228 | 0 | 0.867 | 0 | ||||
| Mistral-7B Instruct-v0.2 | 0.535 | 0 | 0.268 | 0 | 0.883 | 0 | ||||
| Flan-T5 XL | 0.436 | 0 | 0.226 | 0 | 0.875 | 0 | ||||
| Phi-2 | 0.413 | 5 | 0.219 | 10 | 0.858 | 0 | ||||
| Generalist | Flan-T5 XXL | 0.512 | 0 | 0.262 | 0 | 0.885 | 0 | |||
| Llama-2 13B-chat | 0.526 | 0 | 0.273 | 0 | 0.870 | 0 | ||||
| Llama-2 7B-chat | 0.517 | 0 | 0.261 | 0 | 0.868 | 0 | ||||
| Mistral-7B Instruct-v0.2 | 0.542 | 0 | 0.280 | 0 | 0.876 | 0 | ||||
| Flan-T5 XL | 0.463 | 0 | 0.256 | 0 | 0.868 | 0 | ||||
| Phi-2 | 0.479 | 5 | 0.241 | 8 | 0.870 | 0 | ||||
β’
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics βHR@1β and βNDCGβ are detailed in AppendixΒ A. On each task, the best performance is in bold.
Appendix H OOD Evaluation
The following parts discuss the OOD results under different settings.
H.1 Results on Unseen Instructions
TableΒ A17 exhibits the results of evaluating on OOD test sets with unseen instruction. As shown in tableΒ A17, models perform better when training on with diverse instructions. fine-tuned on with diverse instruction enhances the generalizability of models to the unseen instruction.
| Model | Training Instructions | ||||||
| F1* | Macro F1 | Macro F1 | HR@1 | F1 | F | ||
| single | 0.001 | 0.561 | 0.636 | 0.260 | 0.890 | 0.839 | |
| diverse | 0.276 | 0.577 | 0.652 | 0.266 | 0.877 | 0.840 | |
| single | 0.000 | 0.527 | 0.584 | 0.284 | 0.877 | 0.849 | |
| diverse | 0.366 | 0.507 | 0.628 | 0.275 | 0.863 | 0.841 | |
| single | 0.305 | 0.497 | 0.555 | 0.249 | 0.866 | 0.838 | |
| diverse | 0.275 | 0.513 | 0.574 | 0.248 | 0.880 | 0.841 |
β’
In this table, βsingleβ and βdiverseβ indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model tuned over single and diverse instructions on each task is in bold.
H.2 Results on Different Base Models
TableΒ A18 shows the OOD evaluation results of with different base models. With high-quality dataset, models achieve good performance under different base models.
| Model | Base Model | ||||||
| F1* | Macro F1 | Macro F1 | HR@1 | F1 | F | ||
| Flan-T5 XXL | 0.476 | 0.499 | 0.619 | 0.262 | 0.871 | 0.843 | |
| Llama-2 13B-chat | 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 | |
| Llama-2 7B-chat | 0.314 | 0.511 | 0.618 | 0.266 | 0.894 | 0.837 | |
| Mistral-7B Instruct-v0.2 | 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 | |
| Flan-T5 XL | 0.352 | 0.489 | 0.598 | 0.256 | 0.871 | 0.844 | |
| Phi-2 | 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 |
β’
In this table, βBase Modelβ presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
H.3 Results on Models Tuned using All Tasks and Individual Tasks
As demonstrated in TableΒ A19, when generalizing to OOD test, fine-tuned on every single task can better transfer the knowledge to the specific task than fine-tuned on the task combination.
| Model | Training Tasks | ||||||
| F1* | Macro F1 | Macro F1 | HR@1 | F1 | F | ||
| Task-specific | 0.518 | 0.483 | 0.629 | 0.263 | 0.887 | 0.840 | |
| Generalist | 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 | |
| Task-specific | 0.443 | 0.502 | 0.632 | 0.268 | 0.881 | 0.842 | |
| Generalist | 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 | |
| Task-specific | 0.362 | 0.251 | 0.583 | 0.219 | 0.856 | 0.835 | |
| Generalist | 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 |
β’
In this table, βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold.
Appendix I Complete Results for the Analysis of Training Data Size
| Model | Data Size | ||||||||||
| F1* | Macro F1 | F1 | Macro F1 | HR@1 | Accuracy | F1 | NDCG | F1 | F | ||
| 1K | 0.000 | 0.309 | 0.874 | 0.309 | 0.085 | 0.576 | 0.194 | 0.803 | 0.632 | 0.813 | |
| 10K | 0.391 | 0.500 | 0.995 | 0.601 | 0.445 | 0.656 | 0.009 | 0.856 | 0.806 | 0.843 | |
| 47K | 0.544 | 0.549 | 0.995 | 0.618 | 0.506 | 0.675 | 0.424 | 0.869 | 0.854 | 0.844 | |
| 92K | 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 | |
| 1K | 0.046 | 0.327 | 0.982 | 0.550 | 0.318 | 0.633 | 0.156 | 0.817 | 0.792 | 0.828 | |
| 10K | 0.478 | 0.534 | 0.991 | 0.595 | 0.468 | 0.661 | 0.252 | 0.877 | 0.804 | 0.842 | |
| 47K | 0.618 | 0.610 | 0.995 | 0.639 | 0.507 | 0.672 | 0.404 | 0.872 | 0.843 | 0.846 | |
| 92K | 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 | |
| 1K | 0.000 | 0.296 | 0.411 | 0.286 | 0.046 | 0.507 | 0.356 | 0.745 | 0.767 | 0.748 | |
| 10K | 0.223 | 0.330 | 0.987 | 0.510 | 0.287 | 0.636 | 0.000 | 0.838 | 0.772 | 0.840 | |
| 47K | 0.311 | 0.503 | 0.995 | 0.571 | 0.422 | 0.653 | 0.017 | 0.863 | 0.837 | 0.837 | |
| 92K | 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 |
β’
The best performance of , , and over different data sizes is in bold.
| Model | Data Size | ||||||
| F1* | Macro F1 | Macro F1 | HR@1 | F1 | F | ||
| 1K | 0.000 | 0.299 | 0.329 | 0.059 | 0.547 | 0.818 | |
| 10K | 0.389 | 0.464 | 0.602 | 0.227 | 0.833 | 0.842 | |
| 47K | 0.346 | 0.529 | 0.628 | 0.271 | 0.880 | 0.841 | |
| 92K | 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 | |
| 1K | 0.082 | 0.304 | 0.523 | 0.156 | 0.833 | 0.830 | |
| 10K | 0.356 | 0.483 | 0.570 | 0.261 | 0.826 | 0.841 | |
| 47K | 0.401 | 0.529 | 0.650 | 0.267 | 0.891 | 0.846 | |
| 92K | 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 | |
| 1K | 0.000 | 0.278 | 0.278 | 0.052 | 0.801 | 0.758 | |
| 10K | 0.278 | 0.296 | 0.490 | 0.170 | 0.794 | 0.848 | |
| 47K | 0.317 | 0.479 | 0.537 | 0.217 | 0.855 | 0.839 | |
| 92K | 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 |
β’
The best performance of , , and over different data sizes is in bold.
This section exhibits the comprehensive results for analyzing how the training data size influences the performance of models. With the total 92K samples in , we assess data scaling with sizes of 1K, 10K, and 47K. The 1K and 10K samples are collected by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are assembled by randomly selecting 5K samples from each task except for , for which all its 2K training samples are included. As shown in TableΒ A20 and TableΒ A21, performances exhibit an upward trend along with the increase in data size. Notably, the F1 score of the task experiences an initial drop followed by a subsequent rise with the data size increasing. This behavior could be attributed to the imbalanced data, making the models tend to randomly guess when the data size is small. As the data size increases, the models would predict the dominant label, leading to a temporarily low F1 score. With continued data growth, the models acquire sufficient knowledge to generate accurate predictions. In general, the large-scale training data plays a crucial role in developing effective e-commerce LLMs, underscoring the significance of our extensive, comprehensive, and high-quality e-commerce instruction dataset .
Appendix J Case Study
We show the predictions from and the best-performing general-purpose LLM GPT-4 Turbo on each of the 10 tasks. From the following examples, we note that by instruction tuning using our high-quality , could outperform GPT-4 Turbo in distinguishing products (e.g. ), estimating usersβ preferences (e.g. ), and understanding products (e.g. ). For example, as shown in SectionΒ J.2, could accurately predict whether two products are the same given their title, description, manufacturer, and price, while GPT-4 Turbo struggles. In addition, the example in SectionΒ J.5 indicates that compared to GPT-4 Turbo, could better estimate usersβ preferences based on their historical interactions. Moreover, the example in SectionΒ J.6 shows that could correctly understand that DJ light is equivalent to light that will change based on the music. In contrast, without instruction tuning, GPT-4 Turbo may lack this knowledge and fail the example.
J.1 Attribute Value Extraction ()
β’
Instruction:
β
Given the title, description, feature, price, and brand of a product and a set of target attributes, extract the value of each target at tribute from the product information. Output the extracted value and the corresponding source (e.g., title or feature) denoting where the value is extracted.
β
β’
Input:
β
Product title: LibbySue-Airy, Lightweight Knit Winter Scarf in Opulent Earth Tones
β
Product description: The lightweight weave of this soft and airy knit scarf will keep you stylishly ready for any brisk day
β
Product feature: 100% Acrylic, A lightweight weave, Soft and airy knit scarf, Will keep you stylishly ready for any brisk day, Opulent Earth Tones, 78 inches long by 12 inches wide with 7 inch fringe, Package Dimensions: 5.8 x 5.7 x 3.9 inches, Shipping Weight: 4 ounces, ASIN: B00FM6HPAU, Date first listed on Amazon: November 6, 2011, Average Customer Review: 5.0 out of 5 stars 1 customer review
β
Product brand: LibbySue
β
Target attributes: Specialty Style
β
β’
Output:
β
Ground truth: Attribute: specialty style; Value: None, Source: None.
β
: Attribute: specialty style; Value: None, Source: None.
β
GPT-4 Turbo: Attribute: specialty style; Value: airy, lightweight knit, Source: product title.
β
J.2 Product Matching ()
β’
Instruction:
β
Compare the details of two given products to determine if they are identical. Output yes if they are identical or no otherwise.
β
β’
Input:
β
Product 1: title - omniweb 5.0, description - NaN, manufacturer - omni group, price - 29.99
β
Product 2: title - omni web 5.0, description - sure you can use a standard web browser with standard features. but you didnβt choose a standard software experience - you chose the mac. why not try a browser built just for discriminating people with fabulous taste like yourself? omniweb5β¦, manufacturer - NaN, price - 23.99
β
β’
Output:
β
Ground truth: Yes
β
: Yes
β
GPT-4 Turbo: No
β
J.3 Product Relation Prediction ()
β’
Instruction:
β
Given the title of two products, predict if the two products are similar, if the two products will be purchased or viewed together. Answer only from the options.
β
β’
Input:
β
Product 1: Kenable Internal Memory Card Reader for 5.25 CD/DVD Bay With USB Port BLACK
β
Product 2: CORSAIR Carbide 100R Mid-Tower Case
β
β’
Options:
β
A: Users who buy product 1 may also buy product 2.
β
B: Users who view product 1 may also view product 2.
β
C: The product 1 is similar with the product 2.
β
β’
Output:
β
Ground truth: B
β
: B
β
GPT-4 Turbo: A
β
J.4 Sentiment Analysis ()
β’
Instruction:
β
Carefully assess the userβs review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
β
β’
Input:
β
This visor CD holder is a great addition for any car. It folds shut so CDs donβt slide out. It fits well on any visor.
β
β’
Options:
β
A: very positive
β
B: positive
β
C: neutral
β
D: negative
β
E: very negative
β
β’
Output:
β
Ground truth: A
β
: A
β
GPT-4 Turbo: B
β
J.5 Sequential Recommendation ()
β’
Instruction:
β
Given the products the user has purchased in history, rank the items in the listed options and output the item that the user is most likely to purchase next. Answer from one of the options.
β
β’
Input:
β
1st: Bbox A392-10CP Dual 10β Sealed Carpeted Subwoofer Enclosure - Fits 1999-2007 Ford F250/350/450 Crew Cab. Electronics. Car & Vehicle Electronics. b.box.
β
2nd: Smatree Batteries Charger Kit for GoPro Hero 1/2 Digital Camera. Electronics. Camera & Photo. Smatree.
β
3rd: Dolica WT-1003 67-Inch Lightweight Monopod. Electronics. Camera & Photo. Dolica.
β
4th: SunFounder Sidekick Basic Starter Kit w/Breadboard, Jumper wires, Color Led, Resistors, Buzzer For Arduino UNO R3 Mega2560 Mega328 Nano - Including 42 Page Instructions Bookβ¦
β
5th: Winait 5MP Mini 5mp Worlds Smallest Hd Digital Video Camera Spy Camera Video Recorder Hidden Cam DV DVR with 1280 X 960 Resolution. Electronics. Cameraβ¦
β
6th: RoadPro RPPS-220 Platinum Series 12V 3-Pin Plug Fused Replacement CB Power Cord. Electronics. Accessories & Supplies. RoadPro.
β
7th: Transcend USB 3.0 SDHC / SDXC / microSDHC / SDXC Card Reader, TS-RDF5K (Black). Electronics. Computers & Accessories. Transcend.
β
β’
Options:
β
A: Sony Bloggie Live(MHS-TS55) Video Camera with 4x Digital Zoom, 3.0-Inch Touchscreen LCD and WiFi Connectivity (2012 Model). Electronics. Camera & Photo. Sony.
β
B: Gold Tip Expedition Hunter 5575 Dozen Black Shafts. Sports & Outdoors. Sports & Fitness. Gold Tip.
β
C: One Direction Purple Zebra Print Fleece Throw Blanket. Home & Kitchen. Bedding. 1D Media Ltd.
β
D: Absolute DAG15 Dual 15-Inch Angle Ported MDF Enclosure. Electronics. Car & Vehicle Electronics. Absolute.
β
E: Adesso iMouseE1 - Vertical Ergonomic Illuminated Optical 6-Button USB Mouse - Right Hand Orientation. Electronics. Computers & Accessories. Adesso.
β
F: AGPtek AC Power Home Wall Charger Adapter For Microsoft Surface Windows RT Surface 2 Tablet. Electronics. Computers & Accessories. AGPTEK.
β
β¦
β
O: SanDisk Ultra 32GB UHS-I/Class 10 Micro SDHC Memory Card With Adapter - SDSDQUAN-032G-G4A. Electronics. Computers & Accessories. SanDisk.
β
P: UCEC 5.25 Inch Front Panel USB Hub with 2-Port USB 3.0 & 2-Port USB 2.0 & HD Audio Output Port & Microphone Input Port forβ¦
β
Q: GreenLife Soft Grip 11β Ceramic Non-Stick Open Wok, Burgundy. Home & Kitchen. Kitchen & Dining. GreenLife.
β
R: Bulova Frank Lloyd Wright Luxfer Prism Wall Clock, 14β, Bronze. Home & Kitchen. Home Dcor. Bulova.
β
S: Ortopad Girls Eye Patching Reward Posters: 1 Princess Poster, 1 Butterfly Poster. Home & Kitchen. Wall Art. Ortopad.
β
T: HIS H775F1GD Radeon HD 7750 1GB (128bit) GDDR5 Displayport HDMI DVI (HDCP) PCI Express X16 3.0 Graphics Cards. Electronics. Computers & Accessories. HIS.
β
β’
Output:
β
Ground truth: O
β
: O
β
GPT-4 Turbo: A
β
J.6 Multiclass Product Classification ()
β’
Instruction:
β
Determine the relevance between the query and the product title provided, and select your response from one of the available options.
β
β’
Input:
β
Query: dj lights in the car
β
Product title: Sanhezhong Interior Car Lights, LED Car Strip Lights with Waterproof Design, 48 LED Remote Control Car Light Kit, Music Sync Under Dash Car Lighting with Car Charger, DC 12V
β
β’
Options:
β
A: The product is relevant to the query, and satisfies all the query specifications.
β
B: The product is somewhat relevant. It fails to fulfill some aspects of the query but the product can be used as a functional substitute.
β
C: The product does not fulfill the query, but could be used in combination with a product exactly matching the query.
β
D: The product is irrelevant to the query.
β
β’
Output:
β
Ground truth: A
β
: A
β
GPT-4 Turbo: B
β
J.7 Product Substitute Identification ()
β’
Instruction:
β
Assess the relevance of a product to a given query by determining if it can function as a substitute, despite not fully meeting certain aspects of the query. Provide a binary output of yes or no based on this evaluation.
β
β’
Input:
β
Query: iphone 7 plus case otterbox.
β
Product: OtterBox SYMMETRY CLEAR SERIES Case for iPhone 8 Plus & iPhone 7 Plus (ONLY) - Retail Packaging - EASY BREEZY (CLEAR/EASY BREEZY).
β
β’
Output:
β
Ground truth: No
β
: No
β
GPT-4 Turbo: Yes
β
J.8 Query Product Ranking ()
β’
Instruction:
β
Rank the products A, B, C, β¦ based on their relevance to the provided query, and produce a ranked list with the most relevant product positioned at the top of the list.
β
β’
Input:
β
Query: everyone loves raymond dvd complete series
β
Product A: Everybody Loves Raymond: The Complete Series - Seasons 1-9
β
Product B: Perry Mason: The Ninth and Final Season, Vol. 2
β
β’
Output:
β
Ground truth: A, B
β
: A, B
β
GPT-4 Turbo: A
β
J.9 Answerability Prediction ()
β’
Instruction:
β
Given a question and its related document, determine if the question is answerable by analyzing the information in the document. Output yes if the document addresses the question, or no otherwise.
β
β’
Input:
β
Question: if i order 10 of these, are the keys the same for each lock? In other words, can I use any of the keys to open any of the locks? Thanks!
β
Documents:
Document 1: I purchased two of these locks and within two uses with each one I had trouble opening the lock. The third time and I cannot get either opened at all. They are stuck on the tree. Iβll order a different brand and then go cut them off. They keys each turn but they wonβt come off. Wish I had read the reviews.
Document 2: Locks are great my only complaint is I ordered two and I wanted them keyed the same, there was no where in placing the order to specify this and they sent two different keys. Otherwise all is good.
Document 3: I like the Python. Itβs easy to use & I feel it is well made. The problem is I use multiple game cameras & you can not buy the locks keyed alike. My local locksmith who is a Master Lock dealer said can not order them alike. I called Master Locks corporate office-and was told the same. If you need one theyβre great. Iβm forced to look elsewhere.
Document 4: β¦
β
β’
Output:
β
Ground truth: Yes
β
: Yes
β
GPT-4 Turbo: No
β
J.10 Answer Generation ()
β’
Instruction:
β
Answer the given question by extracting information from the supporting document.
β
β’
Input:
β
Question: It says it can support 2 DIMMs and 8GB of RAM. Does this mean I can use a single 8gb stick, or do I have to use 2x4?
β
Documents:
Document 1: Not going to write a full review here, itβs a cheap motherboard, itβs $50, expect to get what you pay for. I bought this from another retailer, but wanted everyone to know that this board does in-fact support 16GB of RAM. I am using this board in a home lab VMware ESXi host, and I needed 16GB at least. Here is what you need to do: (1) Update bios to at least 1303 (here is linky) [β¦] (2) Use RAM on supported RAM list (here is linky)[β¦] I am using 16GB of Corsair VengeanceCorsair Vengeance 16GB (2x8GB) DDR3 1600 MHz (PC3 12800) Desktop Memory (CMZ16GX3M2A1600C10)and it works like a champ.
Document 2: I bought this mobo to replace my ASUS MVA2 HDMI series when upgrading my system. As said in other reviews it can definitely use a 125 watt CPU as I checked the manual before I started my upgrade. Originally this was designed for Phenom 2 X6 cpus which run over 95 watts typically on load.On my rig Iβm using an AMD FX8350, 8gb of Kingston DDR31600 ram (2 sticks of 4gb in each slot dual channel), ASUS AMD Radeon 7750 1gb, Kingston 120gb ssd, 1tb Western Digital hdd, Windows XP 32bit and 7 64 bit, in an Antec desktop case with an upgraded Antec 660 watt power supply.
Document 3: This is the PLUS version and will take a 125W processor (I have a AMD 965 quad core Deneb). It will take 8GB RAM in each of the two slots. Itβs a good board with all the Asus extras on it. Some considerations - the legacy connectors on the back might be useful for you or not. Most people would prefer extra USBβs (or a couple of fast USB). The front panel connector is on the side next to the USB connectors so if you have a short panel connector cord(s) that *might* be an issue for you.
Document 4: Always check the manufacturerβs website for reliable specs. I built a computer using this board and it works perfectly.Reviewer who called this mobo βjunkβ and gave it a β1β was attempting to use a CPU this board isnβt rated for- board rated for up to 95W cpu. No mobo will run with wrong RAM, CPU, inadequate power supply, etc. Double check all components, OS, peripherals, BIOS version and settings, etc. for compatibility.Make sure you know what youβre doing before attempting to build your own computer or you can end up wasting a lot of money.
Document 5: This board is an excellent choice if you are working on a budget and are making a relatively low end system. I used it along with an FX-6100 processor and have had no problems.I would recommend this to anyone building their first system and/or hoping to keep it a cheap build.I would not recommend this as an upgrade but rather a starting point.pros:cheaphigh connectabilitysmallgood SATA port placement.excellent support and manualsimplecons:no USB 3.0 or SATA IIIno IDEonly 4 USB ports on boardlimited ram supportonly PCIe 2.0 x16 @ 8xonly 2 fan connectors (one for cpu and one for case)
Document 6: β¦
β
β’
Output:
β
Ground truth: You can use 8Gb modules, thereβs only 2 DIMMS. Iβm using Corsair DDR 3 16GB kit of 2 8Gb1600Mhz
β
: You can use 2 sticks of 4GB or 1 stick of 8GB.
β
GPT-4 Turbo: You can use a single 8GB stick.
β